### v2

# Efficient Latent Variable Causal Discovery: Combining Score Search and Targeted Testing

Abstract
Learning causal structure from observational data is especially challenging when latent variables or selection bias are present. The Fast Causal Inference (FCI) algorithm addresses this setting but performs exhaustive conditional independence tests across many subsets, often leading to spurious independences, missing or extra edges, and unreliable orientations.
We present a family of score-guided mixed-strategy causal search algorithms that extend this framework. First, we introduce BOSS-FCI and GRaSP-FCI, variants of GFCI (Greedy Fast Causal Inference) that substitute BOSS (Best Order Score Search) or GRaSP (Greedy Relaxations of Sparsest Permutation) for FGES (Fast Greedy Equivalence Search), preserving correctness while trading off scalability and conservativeness. Second, we develop FCI Targeted-Testing (FCIT), a novel hybrid method that replaces exhaustive testing with targeted, score-informed tests guided by BOSS. FCIT guarantees well-formed PAGs and achieves higher precision and efficiency across sample sizes. Finally, we propose a lightweight heuristic, LV-Lite (Latent Variable “Light”), which returns the PAG of the BOSS DAG (Directed Acyclic Graph). Though not strictly sound for latent confounding, LV-Lite often matches FCIT’s accuracy while running substantially faster.
Simulations and real-data analyses show that BOSS-FCI and GRaSP-FCI provide robust baselines, FCIT yields the best balance of precision and reliability, and LV-Lite offers a fast, near-equivalent alternative. Together, these methods demonstrate that targeted and score-guided strategies can dramatically improve the efficiency and correctness of latent-variable causal discovery.
1 Introduction
Learning causal structure from observational data is central in fields such as epidemiology, economics, and biology, where randomized experiments are often infeasible. In many real-world settings, latent variables or selection bias complicate this task, undermining the reliability of traditional discovery methods. The Fast Causal Inference (FCI) algorithm [14] provides a principled framework for such cases by applying conditional independence (CI) tests to subsets of adjacent variables (or Possible-D-SEP sets). While correct in theory, this strategy is fragile in practice: testing many subsets amplifies stochastic variation in $p$-values, leading to spurious independence claims, missing or extra edges, and unstable orientations. We refer to this as the repeated testing problem.
Definition 1 (Repeated Testing Problem).
Let $\mathcal{T}(X{\;\perp\!\!\!\perp\;}Y\mid S)$ denote a statistical CI test at level $\alpha$ for disjoint sets $X,Y,S\subseteq V$. FCI tests many subsets $S\subseteq\text{Adj}(X)\cup\text{Adj}(Y)$ or Possible-D-SEP(X,Y)(X,Y). As $|\mathcal{S}|$ grows, the chance that some $S\in\mathcal{S}$ spuriously satisfies $\mathcal{T}(X{\;\perp\!\!\!\perp\;}Y\mid S)$ increases, inflating Type I errors and producing overly sparse graphs and erroneous orientations.
Beyond sampling variability, another challenge is almost unfaithfulness [18, 15]. Even when the Causal Markov and Faithfulness assumptions hold exactly, small parameter perturbations can produce distributions that are nearly unfaithful—where partial correlations or other dependence measures are close to zero but not exactly zero. In finite samples, such “almost unfaithful” parameterizations lead to unstable independence decisions: small statistical fluctuations can flip dependence judgments, causing orientation errors or spurious adjacencies. As shown by Zhang and Spirtes, [18], this phenomenon can make reliable causal inference difficult even for consistent algorithms, motivating procedures that are robust to weak dependencies or employ stronger assumptions such as $k$-triangle faithfulness [15]. We therefore treat almost unfaithfulness as a complementary concern to the repeated-testing problem—structural rather than statistical in origin, but equally capable of producing unstable PAGs in practice.
Several extensions have sought to mitigate such issues. Order-independent FCI variants [5] stabilize orientations, and score-based methods [^1] Scores in Tetrad currently cover linear Gaussian and linear non-Gaussian models (using BIC-type criteria), multinomial models, mixed multinomial/linear Gaussian models, and, more recently, mixed multinomial/nonlinear post-nonlinear models via truncated basis-function scores [11]. extend causal search to latent-variable models through likelihood-based criteria [16, 4]. Hybrid approaches such as Greedy FCI (GFCI; [10]) combine score-based search with CI testing, but still rely on adjacency-based phases—i.e., during the adjacency-removal stage, testing all subsets of nodes adjacent to a pair—and thus remain vulnerable to repeated testing and almost unfaithfulness. Constraint-based methods, by contrast, have the advantage of being inherently nonparametric and making no distributional assumptions except for those of the selected conditional independence test, which motivates the option of keeping such testing in the algorithm.
Recent advances in score-based search motivate new hybrids that combine the distributional flexibility of constraint-based methods with the efficiency of likelihood-based optimization. Best Order Score Search (BOSS; [2]) provides a scalable, permutation-based alternative to FGES, while GRaSP [9] offers a related variant. Substituting these into the GFCI framework yields BOSS-FCI and GRaSP-FCI—hybrid algorithms that preserve the nonparametric generality of constraint-based search while leveraging score-based initialization to improve stability and scalability. The original GFCI algorithm [10] is sound and complete in the population limit under the Causal Markov and Faithfulness assumptions, returning the correct Partial Ancestral Graph when provided with perfect independence information. However, in finite samples, GFCI and its variants still depend on adjacency-based conditioning and thus remain susceptible to repeated testing and almost unfaithfulness.
Building on these baselines, we propose FCI Targeted-Testing (FCIT), a mixed-strategy algorithm that replaces exhaustive adjacency-based testing with a path-based procedure. Like GFCI, FCIT begins with a score-based completed partially directed acyclic graph (CPDAG)—a representation of the Markov equivalence class of DAGs that share the same adjacencies and unshielded colliders. Score-based algorithms such as BOSS and GRaSP are score-consistent: in the large-sample limit, they recover the correct CPDAG under the Causal Markov and Faithfulness assumptions for the model class (e.g., linear Gaussian, multinomial, or post-nonlinear) used to define the score. FCIT then applies a recursive routine, block_paths_recursively, to identify conditioning sets only when necessary to block active paths between variables, thereby avoiding the over-conditioning that arises in FCI and GFCI from testing all subsets of adjacent variables. Compared with these algorithms, FCIT performs far fewer conditional independence tests, which reduces sampling error and increases the reliability of the overall learned structure. Together with a well-formedness check that enforces the rules of a Partial Ancestral Graph (PAG)—a mixed graph representing the Markov equivalence class of maximal ancestral graphs consistent with possible latent variables and selection bias—this targeted strategy guarantees that every intermediate graph remains structurally consistent, meaning it obeys the ancestral, acyclicity, and maximality constraints required of a valid PAG. The result is greater efficiency, more stable orientations, and structural correctness not ensured by prior FCI variants [17, 8].
Finally, we introduce a deliberately simple heuristic, LV-Lite (Latent Variables “Light”), which outputs the PAG implied by the BOSS DAG without performing any latent-specific conditional independence testing. Though it cannot orient bidirected edges, LV-Lite scales exceptionally well and, in extensive simulation studies, often matches or exceeds the empirical accuracy of more complex algorithms when compared to the true PAG.
Together, these four algorithms—BOSS-FCI, GRaSP-FCI, FCIT, and LV-Lite—form a continuum. The first two establish sound baselines by embedding advanced score-based methods within the GFCI framework. FCIT introduces a principled path-based refinement that improves efficiency and reliability, while LV-Lite represents a heuristic extreme: formally incomplete but practically effective. Our empirical evaluation, spanning both simulation and real data, demonstrates the complementary strengths of these approaches. BOSS-FCI and GRaSP-FCI retain theoretical correctness but are constrained by repeated testing; FCIT balances correctness with improved precision and scalability; and LV-Lite sacrifices correctness guarantees yet achieves strong accuracy with minimal computational cost in simulation.
The remainder of this paper is organized as follows. Section 2 reviews graphical preliminaries. Sections 34 present the four algorithms. Section 5 introduces optimization details, Section 6 outlines theoretical guarantees, and Section 8 reports empirical results. All algorithms are implemented in the open-source Tetrad suite [12], with reproducibility materials available on GitHub.
Contributions.
  • Introduction of BOSS-FCI and GRaSP-FCI, hybrids substituting BOSS and GRaSP for FGES in GFCI.
  • A recursive blocking procedure, block_paths_recursively, that constructs conditioning sets by analyzing paths rather than adjacencies.
  • A targeted-testing algorithm, FCIT, that uses recursive path blocking to maintain well-formed PAGs.
  • A heuristic, LV-Lite, that outputs the PAG of the BOSS DAG, achieving high empirical accuracy at minimal cost.
  • Empirical validation demonstrating accuracy, scalability, and robustness across simulated linear Gaussian datasets and real datasets.
2 Graphical Preliminaries
In this section, we define key terms related to graphical causal models used throughout the paper. We focus on models based on causal directed acyclic graphs (DAGs), which consist of directed edges $X\rightarrow Y$. Such an edge indicates that variable $X$ causes variable $Y$, meaning that interventions on $X$ counterfactually alter $Y$ [14].
A graph $G=(V,E)$ consists of a set of variables $V$ and a set of edges $E$, with at most one edge connecting each vertex pair, where each edge is denoted XYX*{-}*Y. The symbol “*” represents an endpoint, which may be a tail (“-”), an arrowhead (“>>”), or a circle ($\circ$”). Circle endpoints ($\circ$”) indicate that a tail occurs in some members of the corresponding Markov equivalence class, and an arrowhead occurs in others. For example, XYX*{-}*Y could represent $X\circ\!\!\rightarrow Y$, $X\leftarrow Y$, or $X\circ\!\!{-}\!\!\circ Y$. An arrow endpoint $X*\rightarrow Y$ means that $Y$ is not an ancestor of $X$, and a tail endpoint XYX-*Y means that X is an ancestor of Y.
An edge is unoriented if it is of the form $X\circ\!\!{-}\!\!\circ Y$. A directed edge has one arrowhead: $X\rightarrow Y$ or $X\leftarrow Y$. An undirected edge has tails at both ends: XYX-Y. Note that “undirected” refers specifically to tails at both ends, whereas “unoriented” refers to circles at both endpoints. We say $X$ is adjacent to $Y$ in $G$, or $\text{adj}(X,Y)$, if there is an edge XYX*{-}*Y in $G$, where “*” represents a tail (“-”), an arrowhead (“>>”), or a circle ($\circ$”).
A path in $G$ is a possibly empty ordered sequence $\langle X_{1},\dots,X_{n}\rangle$ (written $X_{1}\leadsto X_{n}$) where each pair of vertices adjacent in the sequence is connected by an edge XYX*{-}*Y, where “*” represents a tail (“-”), an arrowhead (“>>”), or a circle ($\circ$”). A directed path in $G$ (written $X_{1}\leadsto X_{n}$) is a path each vertex at position $i\neq j$ in the sequence is connected by an edge $X_{i}\rightarrow X_{i+1}$. A graph is acyclic if it contains no cyclic paths (that is, paths that revisit a node). A node $X$ is an ancestor of $Y$ if either $X=Y$ or else there is a directed path from $X$ to $Y$; $X$ is a parent of $Y$ if $X\rightarrow Y$ is in $G$. A collider on a path is a subpath $\langle A,B,C\rangle$ with $A\rightarrow B\leftarrow C$. It is an unshielded collider if $A$ and $C$ are not adjacent.
We use italic letters for individual variables and boldface for sets of variables. DAGs encode CI statements of the form $X{\;\perp\!\!\!\perp\;}Y\mid\mathbf{S}$ using the d-separation criterion. Specifically, $X$ and $Y$ are d-separated given $\mathbf{S}$ (denoted $\textit{d-sep}(X,Y\mid\mathbf{S})$) if all paths between $X$ and $Y$ are blocked—either because they contain a non-collider in $\mathbf{S}$, or because every collider on the path has no descendant in $\mathbf{S}$.
A graph $G$ is Markov to a distribution $P$ if $\text{d-sep}(X,Y\mid\mathbf{S})$ in $G$ implies that $X$ is independent of $Y$ conditional on $S$, denoted $X{\;\perp\!\!\!\perp\;}Y\mid\mathbf{S}$, in $P$. Conversely, $P$ is faithful to $G$ if $X{\;\perp\!\!\!\perp\;}Y|\mathbf{S}$ implies that $\text{d-sep}(X,Y|\mathbf{S})$.
A Completed Partially Directed Acyclic Graph (CPDAG) represents a Markov equivalence class of DAGs–that is, a set of DAGs all of which imply the same d-separation facts. Markov equivalent DAGs contain the same adjacencies but may differ in orientation. Thus, a CPDAG contains directed edges that appear in all DAGs in the class and undirected edges where orientations differ. Any DAG in the class can be obtained by orienting the undirected edges in a way that maintains acyclicity and preserves the set of unshielded colliders.
A Maximal Ancestral Graph (MAG) is a mixed graph over observed variables encoding the CI facts of a causal DAG with latent and selection variables. It may contain:
  • Directed edges: $X\rightarrow Y$, where $X$ is an ancestor of $Y$;
  • Bidirected edges: $X\leftrightarrow Y$, for latent confounding;
  • Undirected edges: XYX-Y, for selection bias.
MAGs satisfy:
  • Ancestral: No directed or almost directed cycles.
  • Maximality: Every pair of non-adjacent nodes is m-separated by some set.
Selection bias arises when inclusion in the dataset depends on one or more variables in the causal system—typically when data for a variable that is a descendant of multiple causes are incomplete, so that the observed data implicitly condition on that variable. In a causal graph, such conditioning can induce spurious associations among those causes, even when they are otherwise independent. In maximal ancestral graphs (MAGs) and partial ancestral graphs (PAGs), selection bias is represented by undirected edges (XYX{-}Y), indicating that both $X$ and $Y$ are ancestors of a common selection variable. The selection variables in this case are the variables that are being conditioned on in this way.
A Partial Ancestral Graph (PAG) (sometimes referred to as a legal PAG, i.e., simply a PAG) represents a Markov equivalence class of MAGs that imply the same m-separation facts. PAG edges may be:
  • Directed ($\rightarrow$), indicates a causal relationship.
  • Bidirected ($\leftrightarrow$): indicates latent confounding.
  • Undirected (-): indicates selection bias.
Alternatively, PAG edges can contain circle endpoints ($\circ$), indicating a tail occurs in some members of the equivalence class but an arrow in others.
An inducing path between $X$ and $Y$ is a path where every non-endpoint node is a collider and an ancestor of $X$ or $Y$. If such a path exists between non-adjacent nodes, the graph is not maximal.
A MAG or PAG is maximal (following Zhang [17]). if, for every non-adjacent pair $X,Y$, there exists a conditioning set S such that $X$ and $Y$ are m-separated given S. Equivalently, there are no inducing paths between non-adjacent nodes.
An almost-cycle occurs when a graph contains both a bidirected edge $X\leftrightarrow Y$ and a directed path $X\leadsto Y$ (or $Y\leadsto X$), which implies that an arrowhead points to an ancestor, violating acyclicity.
Edges like $X\circ\!\!{-}\!\!\circ Y$ reflect ambiguity at both endpoints. For example, $X\rightarrow Y$ appears in all MAGs in the equivalence class, while $X\leftrightarrow Y$ indicates latent confounding, and $X\circ\!\!{-}\!\!\circ Y$ indicates unresolved ancestral relationships.
A graph G is said to be Markov to the data if the data arise from a probability distribution $P$ over the observed variables and that the conditional independencies implied by G (via d-separation for DAGs or m-separation for MAGs and PAGs) hold in $P$.
A set of variables $\textbf{(}V)$ is causally sufficient if it includes all common causes of observed variables in $\textbf{(}V)$; otherwise, it is causally insufficient. A model is edge-minimal of a certain type (e.g., DAG, CPDAG, PAG) if no model of that type has fewer edges while remaining Markov to the data. Edge minimality in CPDAGs is discussed by Raskutti and Uhler, [13], Lam et al., [9]; here, we extend this to causally insufficient settings.
A path is blocked by a set $\mathbf{Z}$ if:
  • It contains a collider (e.g., $X\rightarrow W\leftarrow Y$) such that $W\notin\mathbf{Z}$ and no descendant of $W$ is in $\mathbf{Z}$;
  • It contains a non-collider node in $\mathbf{Z}$.
A set of nodes $\mathbf{Z}$ blocks a pair of variable sets $\mathbf{X}$ and $\mathbf{Y}$ if every path between any node in $\mathbf{X}$ and any node in $\mathbf{Y}$ is blocked by $\mathbf{Z}$. $X$ and $Y$ are m-separated by $\mathbf{Z}$ if $\neg adj(X,Y)$ and all paths between $X$ and $Y$ are blocked by $\mathbf{Z}$.
An unshielded triple is a sequence $\langle X,Y,Z\rangle$ where $X$ and $Z$ are adjacent to $Y$ but not to each other. If it is oriented as $X*\rightarrow Y\leftarrow*Z$, it is an unshielded collider. These play a central role in the orientation of edges.
A triple $(X,Y,Z)$ forms a definite noncollider at $Y$ if on of the following holds:
  1. 1.
    The triple is unshielded and at least one of the edges XYX*-*Y or ZYZ*-*Y has an arrowhead into $X$ or $Z$, i.e., $Y*\rightarrow X$ or $Y\rightarrow*Z$ exists.
  2. 2.
    The triple is unshielded and both edges have circle endpoints at $Y$, i.e., $X\circ\!\!-\!\!\circ Y\circ\!\!-\!\!\circ Z$, and $X$ and $Z$ are not adjacent.
  3. 3.
    Selection bias is allowed and XYZX-Y-Z.
For any two distinct nodes $A$ and $B$ in a maximal ancestral graph (MAG) $G$, the D-SEP set of $A$ relative to $B$, denoted $\text{D-SEP}(A,B)$, is the set of nodes $V\neq A$ such that there exists a collider path between $A$ and $V$ in which every node is an ancestor of either $A$ or $B$ in $G$. Intuitively, $\text{D-SEP}(A,B)$ contains the variables that may need to be conditioned on to separate $A$ and $B$ when latent confounding is present. In the oracle setting, if $A$ and $B$ are not adjacent in the true MAG, then there exists a subset of $\text{D-SEP}(A,B)$ that d-separates them [14].
Given a PAG $G$, the Possible-D-SEP set of node $A$ includes all nodes $V\neq A$ such that there exists a path $U$ between $A$ and $V$ where each triple $\langle X,Y,Z\rangle$ in $U$ satisfies:
  • $Y$ is a collider, or
  • $Y$ is not a definite non-collider and $\langle X,Y,Z\rangle$ forms a triangle in $G$.
This set may include nonadjacent nodes and is used by FCI and GFCI to search for separating sets in the presence of latent structure [14].
A discriminating path containing nodes in the path $p=\langle X,...,W,V_{xy},Y\rangle$ in a PAG such that:
  • The path contains at least four nodes;
  • $X$ is not adjacent to $Y$, but all intermediate nodes (here, $W$ and $V_{xy}$), and any nodes between $X$ and $W$ are adjacent to $Y$;
  • Every node between $X$ and $Y$ is a collider and a parent of $Y$.
Discriminating paths can imply that a shielded node VxyV{xy} is a collider, helping to orient edges or, in the case of FCIT, to clarify adjacencies in the PAG. Thus, $V_{xy}$ is the node along the path from $x$ to $y$ whose orientation is being discriminated. [^2] The use of discriminating paths is primarily for orientation, though in the case of FCIT in particular, where separating sets are found by looking at longer paths, it is often necessary to know where extra colliders due to such discrimination might be postulated to exist in an evolving PAG. Consider the example in Appendix Section D, for instance.
Key assumptions
  • Causal Markov. For a causally sufficient set of variables, every variable is conditionally independent of its non-descendants given its parents in the given causal DAG.
  • Faithfulness. All CI facts in the distribution follow from the Causal Markov condition applied to the given causal graph.
Note that violations of Faithfulness have Lebesgue measure zero if conditional independence is algebraic, though almost unfaithful violations can be common.
3 Simple Variants of GFCI: BOSS-FCI, GRaSP-FCI, and LV-Lite
The Greedy FCI (GFCI) algorithm [10] is a hybrid causal discovery method combining score-based search with conditional independence (CI) testing. It first estimates a Completed Partially Directed Acyclic Graph (CPDAG) using FGES, then refines it into a Partial Ancestral Graph (PAG) through FCI-style edge removal and orientation rules. This design reduces the search space while remaining robust to latent variables and selection bias.
We introduce three simple variants of GFCI that differ only in the initial scoring method and in whether the subsequent CI-based refinement is applied.
BOSS-FCI.
BOSS-FCI replaces FGES with the Best Order Score Search (BOSS) algorithm [2]. BOSS performs permutation-based score search, offering strong accuracy and scalability—especially for dense graphs. After estimating a CPDAG, the algorithm applies the standard GFCI refinement: adjacency-based edge removal, collider identification, and the FCI orientation rules.
GRaSP-FCI.
GRaSP-FCI substitutes the GRaSP algorithm [9] for FGES in the same framework. GRaSP employs a greedy adjacency search with penalty regularization, emphasizing sparsity and offering a complementary trade-off between accuracy and efficiency. Once the CPDAG is obtained, the usual GFCI refinement steps are applied.
LV-Lite.
LV-Lite [^3] LV-Lite has been referred to as BOSS-POD (BOSS PAG of DAG) elsewhere, but the name “LV-Lite” has stuck, so we defer. is a deliberately simple heuristic. Like BOSS-FCI, it begins with a CPDAG estimated by BOSS, but instead of performing any CI-based refinement, it directly outputs the PAG equivalence class of the resulting DAG. The method therefore captures latent-aware patterns implied by the BOSS structure but cannot remove edges to expose bidirected relationships. Although formally incomplete, LV-Lite is extremely fast and often achieves empirical accuracy comparable to its more principled counterparts.
[…]
Pseudocode.
Appendix A provides pseudocode for the general *-FCI template, where the initial CPDAG procedure $\mathcal{M}$ may be FGES (standard GFCI), BOSS (BOSS-FCI), or GRaSP (GRaSP-FCI). LV-Lite can be viewed as a degenerate instance of this template, omitting edge removal and collider orientation and returning the PAG of the BOSS DAG directly.
Input: PAG $\mathcal{G}$ ; CI oracle $\mathcal{T}$
1
2 Procedure FCI-Backward($\mathcal{G},\mathcal{T}$)
3 repeat
[…]
5
6 until not modified
7 return $\mathcal{G}$
8
9
10 Procedure BackwardStep($\mathcal{G},\mathcal{T}$)
11  for iji*\!-\!*j in $\mathcal{G}$ do
[…]
13        if $i{\;\perp\!\!\!\perp\;}j\mid S$ (via $\mathcal{T}$) then
[…]
16           return true
17
18
19
20 return false
21
Algorithm 1 FCI-Backward (template) with nested BackwardStep. Requirement on the generator PossibleSeparatingSets: it must satisfy the Coverage condition stated below.
e
4 The FCIT Algorithm
Consider an algorithm templated as Algorithm 1, subject to the following conditions:
  • Coverage condition for PossibleSeparatingSets. For every adjacent pair $\{x,y\}$ in the current PAG $\mathcal{G}$, if there exists a set $S^{\star}\subseteq\texttt{PossibleDsep}_{\mathcal{G}}(x,y)$ such that $x{\;\perp\!\!\!\perp\;}y\mid S^{\star}$ (w.r.t. the oracle distribution), then the generator $\tt{PossibleSeparatingSets}(\mathcal{G},x,y)$ must output at least one set $S$ with $x{\;\perp\!\!\!\perp\;}y\mid S$.
  • Initial Superset. The input PAG $\mathcal{G}_{0}$ has an adjacency set that is a superset of the true PAG’s adjacencies.
  • Per-deletion Reorientation and Legality. After each accepted deletion, we apply the complete FCI/Zhang orientation rules to closure and revert if the result is not a legal PAG. This reorientation never introduces new adjacencies.
  • Correct Unshielded Colliders. The input PAG $\mathcal{G}_{0}$ contains all and only the true unshielded colliders implied by the CPDAG. Moreover, after each accepted deletion, any additional unshielded colliders oriented by the FCI/Zhang rules are correct (i.e., consistent with the oracle model).
Correctness sketch (oracle setting). Assume the input PAG satisfies (i) the Initial Superset condition and (ii) the Correct Unshielded Colliders condition. FCIT deletes an edge only when an oracle certifies $X{\;\perp\!\!\!\perp\;}Y\mid S$, and its per-deletion orientation pass is sound, does not create new adjacencies, and only orients true unshielded colliders (illegal states are reverted). By the Coverage property of the separator generator (recursive blocking with PDP forbiddances and $B\setminus D$ trimming), any remaining spurious adjacency XYX*\!-\!*Y is eventually witnessed and removed. Each accepted step strictly reduces the number of edges, so the process terminates. At termination, adjacencies and unshielded colliders are correct; applying the complete Zhang/FCI orientation rules yields the correct (maximally oriented) PAG.
We now instantiate this FCI-Backward template as the FCIT (Fast Causal Inference Targeted-Testing) algorithm, designed to mitigate the repeated-testing problem in FCI-style searches, including other hybrid score-and-test methods. Following GFCI [10], FCIT begins with a score-based procedure that outputs a well-formed CPDAG—e.g., FGES, BOSS, GRaSP, or for small models SP. By default, we use BOSS for its scalability and accuracy.
As noted by Ogarrio et al., [10], score-based CPDAGs can include extra adjacencies or incorrect orientations when latent variables are present, since a CPDAG cannot represent bidirected edges. For instance, the MAG $X\to Y\leftrightarrow Z\leftarrow W$ may yield $X\to Y\to Z\leftarrow W$ (with an extra $X\to Z$) or $X\to Y\leftarrow Z\leftarrow W$ (with an extra $W\to Y$). GFCI corrects these by verifying independence ($X{\;\perp\!\!\!\perp\;}Z$ or $W{\;\perp\!\!\!\perp\;}Y$) and removing the spurious edge, restoring the correct collider structure. FCIT follows the same principle but uses targeted refinement—fewer, higher-value CI tests—while guaranteeing a sound PAG.
Algorithm 2 summarizes the FCIT procedure (see notation in Table 1). The initial CPDAG is converted to a PAG by replacing each edge with $\circ\!$-$\circ$ and transferring the unshielded colliders identified in the CPDAG. This is justified because the final PAG must be a subgraph of the CPDAG’s adjacencies, and collider orientations are the only ones that carry over reliably [10].
The refinement phase employs the block_paths_recursively method (Algorithm 4, Appendix A). Rather than exhaustively test subsets of adjacents or Possible-D-SEP sets, this procedure analyzes individual paths between nodes to construct blocking sets on the fly, selecting at each step a small set of conditional-independence tests to perform, at least one of which is guaranteed to be judged independent precisely in those cases where FCI would find an independence. This strategy substantially reduces the search space, mitigating repeated testing and improving statistical efficiency.
When evaluating an edge xyx*\!-\!*y, FCIT enumerates pre-discriminating paths (PDPs)—paths that could become discriminating for $\{x,y\}$ if the edge xyx*\!-\!*y were removed. [^4] Appendix A.2 gives the efficient algorithm for listing discriminating and pre-discriminating paths. These yield a conservative superset of the true discriminating paths and corresponding $V$-node candidates $\mathsf{V}_{\text{cand},xy}$. Enumerating these supersets allows edges to be processed independently and safely in parallel. For each subset $F\subseteq\mathsf{V}_{\text{cand},xy}$ (including $\emptyset$), FCIT runs recursive blocking to propose candidate separating sets $B\setminus D$ for CI testing. Even when no PDPs are found, FCIT still checks independence with $F=\emptyset$.
Thus, in FCIT, the FCI-Backward generator is instantiated via BOSS initialization and recursive blocking with PDP-based forbiddances and trimming by common adjacents. We prove this separator generator satisfies the Coverage condition. Since FCIT starts from a CPDAG with correct colliders, deletes edges only when $X{\;\perp\!\!\!\perp\;}Y\mid S$ (oracle), and legality-checks each reorientation, every collider added is correct. By Coverage and the Initial Superset invariant, FCIT removes all and only spurious adjacencies and, after the final Zhang pass, returns the correct PAG.
1:Input: data, score method (e.g., BOSS), CI test $\mathcal{T}$
[…]
4: Reset endpoints in $G$ to $\circ\!$-$\circ$
5: Copy known unshielded colliders in $G$ from $C$
6: Apply final FCI orientation rules on $G$, yielding a PAG
[…]
8:while not $\mathit{converged}$ do
9:  $\mathit{converged}\leftarrow\textbf{true}$
10:  for all edges xyx*\!-\!*y in $G$ do
11:   $\mathsf{PDP}_{xy}\leftarrow\textsc{ListPreDiscriminatingPaths}(G,x,y)$
12:   $\mathbf{V}_{\text{cand},xy}\leftarrow\bigcup_{p\in\mathsf{PDP}_{xy}}V(p)$ {superset of true $\mathbf{V}_{xy}$}
13:   $\mathbf{C}_{xy}\leftarrow\mathrm{Adj}_{G}(x)\cap\mathrm{Adj}_{G}(y)$ {common adjacents}
14:   for each $\mathbf{F}\subseteq\mathbf{V}_{\text{cand},xy}$ (including $\emptyset$) $\mathbf{F}$ = forbiddances/not-followed from PDPs}
15:    $\mathbf{B}\leftarrow\textsc{block\_paths\_recursively}(G,x,y,\emptyset,\mathbf{F})$
16:    for each $\mathbf{D}\subseteq(\mathbf{C}_{xy}\cap\mathbf{B})$
17:     if $\exists d\in\mathbf{D}$ with $x\to d\leftarrow y$ then continue {never drop known collider}
18:     $\mathbf{S}\leftarrow\mathbf{B}\setminus\mathbf{D}$
19:     if $x{\;\perp\!\!\!\perp\;}y\mid\mathbf{S}$ (via $\mathcal{T}$) then
20:      Remove xyx*\!-\!*y from $G$
21:      if $G$ is not a legal PAG then revert; continue
22:      $\mathsf{Sep}(\{x,y\})\leftarrow\mathbf{S}$ {cache sepset for this deletion}
23:      Reset endpoints in $G$ to $\circ\!$-$\circ$
24:      Refresh unshielded colliders in $G$ from known list; reapply orientation rules
25:      $\mathit{converged}\leftarrow\textbf{false}$
26:      break both loops
27:  end for
28:end while
29:Output: PAG $G$
Algorithm 2 FCIT (with pre-discriminating paths and recursive blocking)
Table 1: Notation used in FCIT and recursive blocking
Symbol / Name Meaning
$G=(\textbf{V},\textbf{E})$ Current PAG/MAG over observed variables $V$
$x,y$ Distinct query endpoints in $V$
$\mathbf{C}$ Required conditioning set (“containing”)
$\mathbf{F}$ Forbidden nodes (“notFollowing”)
$\mathbf{B}$ Blocking set built by block_paths_recursively
$\mathbf{uc}$ Set of unshielded colliders from CPDAG or updates
$\mathrm{Adj}_{G}(v)$ Adjacency set of node $v$ in $G$
$\mathbf{C}_{xy}$ Common adjacents of $x$ and $y$
$\mathbf{V}_{xy}$ Candidate colliders on discriminating paths w.r.t. $(x,y)$
$\mathbf{Extra}(x,y)$ Certified separating set used to remove xyx*{-}*y
Knowledge Forbidden/required directed edge constraints
$\mathbf{path}$ DFS frontier set to prevent revisits (cycle guard)
$\text{m-sep}(x,y\mid\mathbf{S})$ $x$ and $y$ are m-separated by $\mathbf{S}$ in $G$
5 Optimizing the Final FCI Orientation Rules
A key feature of FCIT is that the final FCI orientation rules [17] are applied repeatedly throughout the refinement process rather than only once at the end. This ensures maximally informative orientations but introduces performance bottlenecks, as several rules are computationally expensive in their standard form. To make FCIT scalable, we optimize the slowest of these rules. Because the same rules are shared across all FCI-style algorithms, the resulting improvements benefit the entire family, including FCI, GFCI, BOSS-FCI, and GRaSP-FCI.
We focus on four rules that dominate runtime in practice. Full details of their optimized implementations are provided in Appendix C.
  • Rule R4 (discriminating paths). Identifying discriminating paths traditionally involves testing subsets of adjacents for nonadjacent pairs. We replace this with the block_paths_recursively procedure, which derives separating sets directly from path structure, avoiding redundant subset enumeration.
  • Rules R5 and R9 (all-circle-endpoint paths). These rules require scanning paths with circle endpoints of potentially unbounded length. We introduce a modified Dijkstra-style shortest-path search that efficiently detects such circle-path structures without exhaustive traversal.
  • Rule R10 (semidirected paths). Detecting uncovered, potentially directed paths can be costly. A path $\langle v_{0},v_{1},\dots,v_{k}\rangle$ is uncovered if no nonconsecutive pair $(v_{i-1},v_{i+1})$ is adjacent in $G$. We implement a memoized depth-first search that enforces uncoveredness and directional constraints exactly while pruning redundant traversals.
These optimizations substantially reduce both the number of CI tests and the number of path enumerations required. In FCIT—where orientation rules are invoked repeatedly—these savings compound to large runtime improvements. The same implementations can be applied directly to other FCI-style algorithms, improving efficiency without altering correctness. [^5] For consistency with the original algorithm definition, for FCI we use the original adjacency-based discriminating path rule, though for the other algorithms we use the recursive one.
6 Theoretical Guarantees
We outline the guarantees enjoyed by our introduced algorithms, including FCIT, under the standard Causal Markov and Faithfulness assumptions (Section 2), assuming access to an oracle conditional independence (CI) test. Finite-sample behavior is examined in Section 8.
We first summarize the correctness status of the simpler variants introduced in Section 3, then turn to FCIT itself. For particular choices of score and/or test, the procedure will make any additional assumptions which chosen score and/or test themselves make, though correctness of the procedures themselves, given non-parametric tests or scores, is itself non-parametric.
6.1 Correctness of BOSS-FCI and GRaSP-FCI
BOSS-FCI and GRaSP-FCI inherit the correctness guarantees of the GFCI algorithm [10]. In its original form, GFCI included an additional Possible-D-SEP search step that ensures oracle correctness. Although this step is often omitted in practice, and no counterexamples have been observed even under a $d$-separation oracle, a formal proof of its redundancy remains open. With that step included, BOSS-FCI and GRaSP-FCI are correct for the same reason as GFCI, differing only in the score-based method used for the initial CPDAG.
6.2 Correctness of LV-Lite
By contrast, LV-Lite lacks formal correctness guarantees. It reports only the PAG equivalence class of the DAG returned by BOSS and therefore cannot remove edges to expose latent confounding or orient bidirected edges. Nevertheless, it always produces a structurally well-formed PAG and, as shown in Section 8, performs well on adjacency and orientation accuracy measures.
To provide intuition for why LV-Lite, while incorrect, nevertheless achieves strong performance on PAG accuracy measures, we note the following considerations. It is unsurprising that LV-Lite performs well on arrowhead precision for common adjacencies (AHPC), since it should in principle produce a subset of the arrowheads that exist over the adjacencies shared with the true PAG. Similarly, adjacency precision (AP) tends to be high: by GFCI theory, LV-Lite should add only a small fraction of “extra” adjacencies, primarily to cover colliders present in the true PAG. Arrowhead recall (AHR) is also expected to be good if the hidden arrowheads represent only a small proportion of those in the true PAG.
Taken together with the high accuracy of BOSS for producing estimates of CPDAGs, these and other such effects explain why LV-Lite generally yields graphs that are very close to the true PAG—especially in large models—despite its lack of formal guarantees. [^6] We did not pursue, but plan to in future work, estimating effects of direct intervention.
6.3 Correctness of FCIT
FCIT introduces a recursive blocking procedure that replaces adjacency-only conditioning and integrates discriminating paths directly into edge removal. The following results summarize its guarantees in the oracle setting; detailed proofs appear in Appendix B.2 and B.3.
Theorem 1 (Soundness of Recursive Blocking).
Let $x\neq y$ be distinct measured vertices in a graph $G$, where $G$ is a PAG (or, more generally, a MAG or DAG) interpreted under $m$-separation. Let $\mathit{block\_paths\_recursively}(x,y,C,\mathbf{F})$ denote Algorithm 4. If the algorithm halts and returns a blocking set $\mathbf{B}$, then either $x$ and $y$ are adjacent, or every path $p$ from $x$ to $y$ in $G$ is blocked by $\mathbf{B}$:
$adj(x,y)\;\;\vee\;\;\forall p.\;\operatorname{Path}(x,y,p)\Rightarrow\operatorname{Blocked}(p,\mathbf{B}).$
Corollary 1 (Adjacency Case).
If $\mathit{block\_paths\_recursively}(x,y,C,\mathbf{F})$ halts and returns a non-null $\mathbf{B}$ with $adj(x,y)$ in a MAG or PAG $G$, then $\mathbf{B}$ blocks all xxyy paths except the direct single-edge path:
$adj(x,y)\wedge\mathbf{B}\neq\textsc{Null}\;\Longrightarrow\;\forall p.\;\operatorname{Path}(x,y,p)\wedge|p|\!\geq\!2\Rightarrow\operatorname{Blocked}(p,\mathbf{B}).$
Intuition. When $x$ and $y$ are adjacent in a MAG, the recursive procedure still verifies that every indirect route $x,b,\ldots,y$ through any neighbor $b\!\neq\!y$ is $m$-blocked by $\mathbf{B}$. If the algorithm returns a non-null $\mathbf{B}$, all non-inducing paths are closed, leaving only the direct xyx*\!-\!*y link potentially open in the corresponding MAG or PAG representation.
6.4 Edge-Minimality of the Returned PAG
Lemma 1 (RB completeness over $\bm{\mathcal{S}}$).
If there exists $\mathbf{S}\in\bm{\mathcal{S}}(X,Y)$ such that $X{\;\perp\!\!\!\perp\;}Y\mid\mathbf{S}$ (in the oracle PAG), then $\mathit{block\_paths\_recursively}(X,Y,C,\textbf{F})$ halts and returns some blocking set $\mathbf{B}\subseteq\bm{\mathcal{S}}(X,Y)=\mathrm{Adj}(X)\cup\mathrm{Adj}(Y)\cup\mathrm{Possible\text{-}D\text{-}SEP}(X,Y)$ with $X{\;\perp\!\!\!\perp\;}Y\mid\mathbf{B}$.
Intuition. The recursive blocking procedure explores all open paths between $X$ and $Y$ and adds noncolliders that lie on those paths until every path is $m$-blocked. Every node it ever considers conditioning on lies on a path between $X$ and $Y$ whose internal vertices are either colliders or ancestors of colliders—exactly the definition of Possible-D-SEP. Hence, any node that recursive blocking may add is already contained in $\bm{\mathcal{S}}(X,Y)$. If some separating set $\mathbf{S}\in\bm{\mathcal{S}}(X,Y)$ exists, recursive blocking will therefore discover a (possibly different) subset $\mathbf{B}\subseteq\bm{\mathcal{S}}(X,Y)$ that also blocks all XXYY paths.
Theorem 2 (Edge-Minimality).
Let $G$ be the true PAG and $G^{\prime}$ the PAG returned by FCIT. If an edge XYX*\!-\!*Y remains in $G^{\prime}$, then there exists no conditioning set $\mathbf{S}$ from the standard FCI search space $\bm{\mathcal{S}}(X,Y)=\mathrm{Adj}(X)\cup\mathrm{Adj}(Y)\cup\mathrm{Possible\text{-}D\text{-}SEP}(X,Y)$ that $d$-separates $X$ and $Y$.
Proof sketch. If such a $\mathbf{S}$ existed, Lemma 1 implies that $\mathit{block\_paths\_recursively}(X,Y,C,\textbf{F})$ would find some blocking set $\mathbf{B}\subseteq\bm{\mathcal{S}}(X,Y)$ with $X{\;\perp\!\!\!\perp\;}Y\mid\mathbf{B}$. FCIT would then test this independence and remove the edge. Since FCIT removes no other edges, any adjacency that remains corresponds to a pair $(X,Y)$ for which no valid separating set exists within $\bm{\mathcal{S}}(X,Y)$. Hence, $G^{\prime}$ is edge-minimal.
6.5 Soundness of Orientations
Theorem 3 (Orientation Soundness).
Let $G$ be the true PAG and $G^{\prime}$ the PAG returned by FCIT. Every oriented edge in $G^{\prime}$ (arrowhead or tail) appears with the same orientation in $G$.
Proof sketch. FCIT uses exactly the arrow- and tail-complete rule set R1–R10 of Zhang, [17]. Given correct adjacencies and unshielded colliders, Spirtes and Zhang prove each rule sound; hence all orientations in $G^{\prime}$ are valid.
6.6 Optional Completeness of Orientations
Theorem 4 (Orientation Completeness).
FCIT produces a tail- and arrow-complete PAG (i.e., every orientation entailed by Markov+Faithfulness appears) iff the full rule set R1–R10 is applied to closure. If any rule is disabled or the process halts early, the resulting PAG may be partially oriented, but all stated orientations remain sound by Theorem 3.
6.7 Summary Corollary
Corollary 2.
Under Causal Markov and Faithfulness, the PAG $G^{\prime}$ returned by FCIT (i) is Markov to the observed distribution, (ii) is edge-minimal with respect to the standard FCI conditioning space, and (iii) contains only sound orientations. If all rules R1–R10 are applied to closure, $G^{\prime}$ is additionally tail- and arrow-complete.
7 Well-formedness of PAGs and Markov
One observes that FCIT and LV-Lite always return graphs that are well-formed, in the sense that they satisfy strictly the definition of a PAG. (For a simulation comparison of how well these various algorithms recover PAGs, consider Figure 12 and related figures in the Appendix.) We can determine this independently by checking whether the identified Zhang MAG calculated from estimated graph [17] satisfies the definition of a MAG and whether when converted back to a PAG the orientations of the estimated PAG are recovered precisely. That is, we check the following
  • No Almost-Cycles or Directed Cycles. We check in the Zhang MAG whether any bidirected edge $A\leftrightarrow B$ has a directed path $A\leadsto B$ or $B\leadsto A$ or whether there are any directed cycles from nodes to themselves. If not, the final graph cannot exhibit an almost-cycle or a directed cycle in the induced MAG.
  • Maximality. Next, we check to make sure the graph does not contain any inducing paths from nodes $A$ to $B$ in the graph while $\neg adj(A,B)$, in the Zhang MAG of the estimated graph. If not, then the MAG is maximal.
  • Orientation Rule Compliance. Finally, we check whether converting the Zhang MAG back into the equivalence class PAG it is a member of yields a graph that has the same endpoint markings as the original estimated graph. If not, then the graph lies somewhere between a MAG and a PAG.
These three conditions are individually necessary and jointly sufficient for a singly connected graph over measured variables alone to be a structurally well-formed PAG; we can therefore verify well-formedness by checking these conditions directly. [^7] The Tetrad suite of software contains such a check. By construction, FCIT and LV-Lite always return structurally well-formed PAGs, whereas empirically we find that other FCI-style procedures do not always do so.
In particular:
  • FCIT. By construction, FCIT checks every proposed update for structural validity, ensuring that the algorithm always moves among well-formed PAGs. Consequently, its output is guaranteed to be structural well-formed regardless of unfaithfulness or finite-sample noise.
  • LV-Lite. LV-Lite trivially returns a well-formed PAG by construction, since it outputs the PAG equivalence class of the BOSS output, which is guaranteed to be a DAG (as BOSS is an algorithm that infers a DAG from a particular permutation of the variables).
  • BOSS-FCI GRaSP-FCI, and GFCI. Under unfaithfulness, these algorithms can return malformed graphs in practice.
  • FCI. Under unfaithfulness, FCI can also produce a graph that is not a well-formed PAG.
Note that each of these algorithms in Oracle mode is guaranteed to produce a well-formed PAG.
We next compare FCIT with FCI, GFCI, BOSS-FCI, GRaSP-FCI, and the heuristic LV-Lite using simulations implemented in Tetrad [12], focusing on accuracy and efficiency in the presence of latent variables.
8 Empirical Evaluation
8.1 Simulation Results
8.1.1 Simulation Setup
We simulated data from linear Gaussian structural equation models generated from random forward DAGs with latent variables. [^8] As suggested in an earlier footnote, this is a modeling choice in Tetrad. Scores and tests are also available for the multinomial case, the mixed multinomial/Gaussian case, and the mixed multinomial/post-nonlinear case (see [11]. Also, it should be noted that linear Gaussian scores and tests also work well in the linear non-Gaussian regime [2]). Tests are also available for the general non-parametric case, though not scores currently; these are less scalable. The data were produced using a standard linear Gaussian simulation method (SemSimulation in Tetrad) with the following design:
  • Graph size and density. $20$ measured variables with average degree $2$, 44, or $6$. For larger regimes we used $100$ measured variables (avg. deg. $6$, $10$ latents) and a single stress test with $200$ variables (avg. deg. $6$, 2020 latents).
  • Latent structure. Each $20$-node run included 0, 44, or $8$ latent variables in addition to the measured ones. Latent variables were not saved out but contributed to inducing dependence among observed variables.
  • Coefficients. Edge coefficients were drawn uniformly from [1,1][-1,1], allowing both positive and negative effects. (Note that we do not use an interval about zero.)
  • Variances. Independent noise variances were sampled from $U(1,3)$; no additional measurement noise was added.
  • Sample sizes. $N=200,500,1000,5000$ for the $20$-node grids. The $100$- and $200$-node settings used $N=1000$.
  • Replicates. Each condition was repeated $20$ times with randomized column order.
Throughout, the score was the linear Gaussian BIC score with penalty discount $2.0$, and the conditional independence test was Fisher’s $Z$-test at $\alpha=0.01$. These parameters did not vary across conditions. Graphs were not required to be connected, and maximum degree limits were set sufficiently high to avoid constraining the generated structures.
8.1.2 Evaluation Metrics
We report standard adjacency and orientation measures, along with three path-aware precision statistics that compare estimated endpoints to directed reachability in the true graph.
  • Adjacency: precision and recall for adjacency recovery (AP, AR).
  • Arrowheads and tails: precision and recall for arrow correctness (AHP, AHR); precision and recall for arrow correctness for adjacencies in common with the true graph (AHPC, AHRC).
  • Arrow Path Precision (no back-reach): proportion of estimated edges $X*\!\to Y$ such that there is no directed path $Y\leadsto X$ in the true DAG (*->-Prec)
  • Tail Path Precision (forward reach): proportion of estimated edges $X\to Y$ such that there is a directed path $X\leadsto Y$ in the true DAG. (–>-Prec)
  • Bidirected Latent Precision: proportion of estimated $X\leftrightarrow Y$ edges for which a latent variable $L$ exists with $X\leftarrow L\rightarrow Y$ in the generating DAG (<->-Lat-Prec)
  • Well-formed PAGs: proportion of discovered graphs that satisfy the structural definition of a valid PAG.
For AP, AR, AHP, AHR, AHPC, and AHRC, estimated graphs are compared to the true PAG; for *->Prec, –>-Prec, and <->-Lat-Prec, graphs are compared to the true DAG (with latents).
We also report runtime (E-CPU) to assess computational efficiency.
8.1.3 20 nodes.
Results for the 20-node case with average degree 4 are shown in Figures 113. Each plotted point is the average of 20 runs per condition, with sample sizes $N\in\{200,500,1000,5000\}$ (log-scaled on the x-axis) and averaged across 0, 4, and 8 latent common causes. All metrics here compare each estimated graph to the true PAG.
All methods achieve very high adjacency precision. FCIT and LV-Lite consistently lead most orientation-precision metrics (AP/AHP, AHPC, arrow-path precision) and always return valid PAGs. Tail-path precision is highest for BOSS-FCI and GFCI at large $N$, with FCIT close behind. FCIT attains the best bidirected latent-path precision at large samples, while LV-Lite—by design—does not orient bidirected edges. In runtime, LV-Lite (and typically BOSS-FCI) are fastest; GFCI is slowest; FCIT and GRaSP-FCI fall in between. Results for average degrees 2 and 6 (see Appendix) show qualitatively similar patterns, with denser graphs highlighting the tail-path strengths of BOSS-FCI/GFCI and the robustness of FCIT/LV-Lite on PAG validity.
8.1.4 100 nodes.
[…]
At this larger scale, LV-Lite and FCIT both maintain exceptionally high arrow-path precision ($\approx 0.99$) and consistently produce structurally valid PAGs, though their tail-path precision is somewhat lower ($\approx 0.93$). In contrast, BOSS-FCI and GRaSP-FCI display the opposite trade-off—achieving very high tail-path precision ($\approx 0.97$) but lower arrow-path precision ($\approx 0.89$). Their PAG validity, however, drops to roughly 0.10–0.25. GFCI performs comparably on both arrow and tail paths but with reduced latent-edge precision and moderate PAG validity (0.20), whereas FCI shows the weakest precision scores overall but remains by far the fastest method. Bidirected-edge precision remains low for all methods at this scale, suggesting that in large models such edges should be interpreted cautiously or reserved for qualitative guidance. Overall, these results mirror the smaller-graph findings: FCIT offers the best balance of accuracy, scalability, and structural validity, while LV-Lite remains the most efficient and structurally stable baseline.
8.1.5 Takeaway.
Across sparsity levels, FCIT and LV-Lite remain the most accurate and stable overall. Both consistently produce valid PAGs, sustain near-perfect arrow- and tail-path precision, and recover adjacencies reliably even under latent confounding. LV-Lite matches FCIT’s orientation accuracy while running substantially faster, making it a strong baseline when latent structure need not be modeled explicitly. BOSS-FCI and GRaSP-FCI maintain high directional precision but occasionally yield malformed PAGs and incur heavier computational cost at scale. GFCI continues to perform reasonably well on directed edges yet shows moderate PAG validity and weaker latent-edge precision, while standard FCI remains the fastest but the least accurate overall. In sum, FCIT offers the best balance of correctness, scalability, and structural validity among algorithms that consistently return legal PAGs across graph densities.
Figure 1: Adjacency Precision (AP) for 20-node graphs with average degree 4. All algorithms maintain very high adjacency precision across sample sizes. LV-Lite trails slightly. LV-Lite is slightly lower than the others but remains acceptable.
Figure 2: Adjacency Recall (AR) for 20-node graphs with average degree 4. LV-Lite and FCIT achieve the highest recall across all sample sizes, reflecting their ability to recover true adjacencies even in the presence of latent confounding. BOSS-FCI, GRaSP-FCI, GFCI, and FCI improve steadily with $N$. Results are averaged over graphs with 0, 4, and 8 latent common causes.
Figure 3: Arrowhead Precision (AHP) for 20-node graphs with average degree 4. LV-Lite and FCIT maintain the highest precision, followed by BOSS-FCI and GRaSP-FCI. FCI remains lowest throughout, while GFCI improves moderately at large NN. Results are averaged over graphs with 0, 4, and 8 latent common causes.
Figure 4: Arrowhead Recall (AHR) for 20-node graphs with average degree 4. FCIT and LV-Lite show the strongest growth in recall with increasing NN, approaching 0.6 at large NN. BOSS-FCI, GFCI, and GRaSP-FCI track closely, while GFCI remains lower overall. Results are averaged over graphs with 0, 4, and 8 latent common causes.
Figure 5: Arrowhead Precision for Common Adjacencies (AHPC) for 20-node graphs with average degree 4. LV-Lite and FCIT achieve the highest arrowhead precision, with LV-Lite reaching 0.92 at large NN. BOSS-FCI and GRaSP-FCI show similar trends, while GFCI and especially FCI remain lower. Results are averaged over graphs with 0, 4, and 8 latent common causes.
Figure 6: Arrowhead Recall for Common Adjacencies (AHRC) for 20-node graphs with average degree 4. All algorithms improve steadily with increasing NN, with FCI, GRaSP-FCI, and BOSS-FCI converging near 0.9. FCIT levels off slightly below these, while LV-Lite stabilizes around 0.85. GFCI performs comparably to the best methods at large samples. Results are averaged over graphs with 0, 4, and 8 latent common causes.
Figure 7: Tail Precision (TP) for 20-node graphs with average degree 4. GFCI, LV-Lite, and LV-Lite lead here, with other algorihtms lagging behind. This metric measures tail precision relative to the true PAG; for a comparison to the directed paths in the true DAG, see Tail Path Precision.
Figure 8: Tail Recall (TR) for 20-node graphs with average degree 4. LV-Lite and FCIT achieve the highest recall across all sample sizes, reflecting their ability to recover correct tail orientations even in the presence of latent confounding. The other algorithms are grouped substantially lower. Results are averaged over graphs with 0, 4, and 8 latent common causes.
Figure 9: Arrow Path Precision for 20-node graphs with average degree 4. LV-Lite and FCIT again lead in directional precision, followed closely by BOSS-FCI, GRaSP-FCI, and GFCI. FCI trails across sample sizes. Results are averaged over graphs with 0, 4, and 8 latent common causes.
Figure 10: Tail Path Precision for 20-node graphs with average degree 4. The algorithms are quite competitive on this measure, with FCI lagging somewhat behind. Results are averaged over graphs with 0, 4, and 8 latent common causes.
Figure 11: Bidirected Latent Path Precision for 20-node graphs with average degree 4. Averaged over the number of latent variables, bidirected edge precision is somewhat variable across methods. FCIT is the most consistent across sample sizes. Results are averaged over graphs with 0, 4, and 8 latent common causes.
Figure 12: Proportion of well-formed PAGs for 20-node graphs with average degree 4. As theoretically required, LV-Lite and FCIT achieve perfect PAG validity, whereas other algorithms fall off—sometimes dramatically. Results are averaged over graphs with 0, 4, and 8 latent common causes.
Figure 13: Runtime (CPU milliseconds) for 20-node graphs with average degree 4. All algorithms are quite fast in this regime, with the exception of GFCI, which slows at N=5000N{=}5000. LV-Lite lags behind LV-Lite. Results are averaged over graphs with 0, 4, and 8 latent common causes.
Table 2: Comparison of algorithms on precision and PAG metrics.
Algorithm *->-Prec –>-Prec <->-Lat-Prec E-CPU PAG
LV-Lite 0.9939 0.9320 * 21850.5654 1.0000
FCIT 0.9937 0.9319 0.0000 24430.5061 1.0000
BOSS-FCI 0.8971 0.9763 0.0961 23183.7960 0.1000
GRaSP-FCI 0.8889 0.9693 0.0966 92879.1827 0.2500
GFCI 0.7861 0.8812 0.0604 54695.2840 0.2000
FCI 0.6519 0.8200 0.0442 1253.2596 0.0500
8.2 Real-Data Analysis
We analyzed the Algerian Forest Fire dataset from the UCI Machine Learning Repository [1]. Because we have not discussed nonlinearity or determinism, we removed the six deterministic Canadian Forest Fire System indices—BUI (Build-Up Index), DC (Drought Code), DMC (Duff Moisture Code), FFMC (Fine Fuel Moisture Code), and FWI (Fire Weather Index). We also excluded Day (cyclical variable) and Year (constant at 2012 and therefore non-causal). This left the variables Region (binary), Month, Relative Humidity (RH), Rain, Temperature, Wind Speed (Ws), and Fire (binary occurrence of fire).
Because the dataset mixes discrete and continuous variables, we used the Degenerate Gaussian score and test [2], with penalty discount $1.0$ and $\alpha=0.05$. This method models discrete variables as multinomial and continuous variables as linearly related, enabling hybrid analysis under a unified likelihood framework.
The resulting FCIT model (Figure 14) closely follows the initial BOSS CPDAG while clarifying uncertain endpoints in the inferred PAG. Notably, Fire has two direct causes—Temperature and Relative Humidity—with Rain as a possible third, a pattern that is meteorologically plausible. Month and Region are correctly identified as exogenous, and the remaining edges are sensible.
For comparison, LV-Lite returns a PAG nearly identical to the BOSS CPDAG, without introducing bidirected edges. BOSS-FCI and GRaSP-FCI produces similar but slightly denser graphs.
In Ramsey and Andrews, [11], we conducted a complementary analysis using ChatGPT-4 to provide “expert” judgments of edge and orientation correctness for this dataset. This included models where the degenerate Gaussian score was applied without the Canadian Fire System variables (i.e., the same configuration analyzed here), in cases where no domain expert was available. In that study, the FCIT graph was judged to offer the most accurate overall summary, whereas alternative models exhibited spurious adjacencies or misoriented edges. FCIT preserved the accuracy of the BOSS initialization while refining orientations in a manner consistent with PAG semantics, highlighting its advantage—when paired with the degenerate Gaussian test and score—for real-world data containing mixtures of continuous and discrete variables.
Beyond our own benchmark datasets, FCIT and the related nonlinear basis-function test/score [11] have recently been applied to astrophysical data at scale. [6] applied FCIT to a highly nonlinear, non-Gaussian sample of roughly $5\times 10^{5}$ low-redshift galaxies from the NASA Sloan Atlas. Using the Basis-Function LRT for independence testing and Basis-Function BIC for scoring, they recovered a causal graph revealing hierarchical, mass-driven relations among galaxy properties while distinguishing them from observational selection effects, which in this note was judged to be quite plausible, recovering known relationships among the studied variables. This study, submitted to MNRAS, demonstrates that FCIT with an appropriate test and score scales to genuinely astronomical sample sizes and provides interpretable and plausible causal hypotheses for physical mechanisms—marking the first published scientific deployment of FCIT.
Figure 14: FCIT result for Algerian Forest Fire data using the Degenerate Gaussian score and test, excluding deterministic indices.
9 Discussion and Conclusion
We have introduced a family of score-guided, mixed-strategy algorithms for causal discovery under latent variables and selection bias. BOSS-FCI and GRaSP-FCI extend the GFCI framework by substituting BOSS or GRaSP for FGES in the initial CPDAG search. LV-Lite provides a simple yet effective heuristic alternative that returns the PAG equivalence class of the BOSS DAG, scaling extremely well but lacking the ability to remove edges indicative of latent confounding. Finally, FCIT introduces a targeted-testing strategy that addresses a key statistical limitation of FCI-style methods.
A central innovation of FCIT is its integration of recursive path blocking into edge removal, in conjunction with discriminating paths, allowing separating sets to be identified using full path structure rather than adjacency subsets alone. Each update is checked for structural validity, guaranteeing that FCIT always produces a well-formed PAG. Our theoretical results establish the soundness of recursive blocking, the edge-minimality of the returned PAG, and the soundness of all orientations in the large-sample limit. Empirically, FCIT achieves consistently high adjacency and orientation precision, scales effectively to large models, and outperforms existing FCI-style methods in both accuracy and efficiency. Although not formally correct, LV-Lite remains a remarkably fast and accurate heuristic for very large problems.
These contributions open several directions for future work. First, automated strategies for jointly tuning score penalties and test thresholds could adaptively balance Type I and Type II errors in practice. Second, extending FCIT to dynamic or time-series data is a natural next step, as are variants that incorporate interventional or mixed observational–interventional data. Finally, while FCIT infers PAGs that reflect latent confounding, it does not explicitly model the latent variables themselves; integrating latent-variable identification within the same framework remains an important challenge.
Finally, we would be remiss not to note an important opportunity for future software development relevant to all algorithms in this paper that begin with a BOSS initialization—LV-Lite, BOSS-FCI, and FCIT—for the linear Gaussian or linear non-Gaussian cases. Our implementations have focused on Java for portability and integration with the Tetrad platform, which already provides strong performance advantages over many alternatives. However, a recent C-language implementation, CBOSS, available in the causal-get GitHub package (https://github.com/bja43/causal-get, due to Bryan Andrews), achieves substantial speed improvements. [^9] The speed improvements are largely due to low-level linear algebra optimizations for calculating likelihoods. The practical extension to the linear non-Gaussian case is benchmarked in Andrews et al., [3]. For problems with 100 nodes and average degree 10—the “large” case profiled above—CBOSS completes in under a second with performance identical to our Java-based BOSS and scales efficiently to much larger problems. This opens the possibility of integrating CBOSS as an external module for the BOSS phase of our algorithms. In Java, such integration poses some engineering challenges, but a hybrid approach—running the BOSS stage in C and importing the resulting graph into Java for subsequent processing—would be straightforward. Alternatively, the remaining steps of our BOSS-based algorithms could themselves be reimplemented in C. Either direction represents a practical and worthwhile programming goal for future work.
In summary, FCIT provides a practical and statistically efficient alternative to existing FCI-style algorithms, consistently yielding well-formed PAGs with improved accuracy and scalability. Together with BOSS-FCI, GRaSP-FCI, and the heuristic LV-Lite, it expands the toolkit for causal discovery under latent variables—bridging the gap between correctness, scalability, and practical usability.
Acknowledgments and Disclosure of Funding
Thanks to Clark Glymour for helpful comments.
JR was supported by the US Department of Defense under Contract Number FA8702-15-D-0002 with Carnegie Mellon University for the operation of the Software Engineering Institute. The content of this paper is solely the responsibility of the authors and does not necessarily represent the official views of these funding agencies. BA was supported by the US National Institutes of Health under the Comorbidity: Substance Use Disorders and Other Psychiatric Conditions Training Program T32DA037183. The authors have no conflict of interest to report. PS was supposed by NIH Award Number: 1043252, “Interpretable graphical models for large multi-model COPD data” and NSF award number 2229881, “AI Institute for Societal Decision Making.”
Appendix A Algorithm Details
A.1 GFCI Variants, Find Path to Target, Recursive Blocking
The structure of GFCI, BOSS-FCI and GRaSP-FCI is shown in Algorithm 5, which we refer to using the templated name, “Star-FCI”. Note that we do not include the possible d-sep step in this code which was included in the published version of GFCI, following our conjecture that it is not needed.
For FCIT, The crucial algorithm find_path_to_target is given in Algorithm 6, along with helper methods in Algorithm 3. This is used in the algorithm, block_paths_recursively, Algorithm 4, to construct sets B on the fly to block all paths between variables $X$ and $Y$. This uses the notion of reachability given in [7]. The optimization of the discriminating path rule R4 is given in Algorithm 7. Section A.2 gives the procedure we use to list discriminating and pre-discriminating paths.
As an intuition, the recursive blocking procedure underlying FCIT is closely related to the standard depth-first search algorithm for determining d-separation between variables $X$ and $Y$ given a conditioning set $\mathbf{Z}$ [7]. Its main innovation is to construct the conditioning set dynamically: as the algorithm explores paths from $X$ to $Y$, it incrementally builds a blocking set $\mathbf{B}$ by adding nodes that close open trails—“in a clever way.” [^10] We borrow this turn of phrase from Zhang, [17]. When the procedure halts and $X$ and $Y$ are not adjacent, it returns a final set $\mathbf{B_{final}}$ such that $\text{d-sep}(X,Y\mid\mathbf{B}_{final})$, or Null if no such separating set exists. The same logic applies to DAGs (directed acyclic graphs), CPDAGs (completed partially directed acyclic graphs), MAGs (mixed ancestral graphs), and PAGs (partial ancestral graphs). If $X$ and $Y$ are adjacent, the procedure instead returns a set that blocks all non-inducing paths between them—a property that FCIT directly exploits.
1Goal: Determine which adjacent nodes $c$ of $b$ can be traversed, given a prior node $a$ and a conditioning set $\mathbf{B}$.
2 Procedure: $\texttt{getReachableNodes}(\textit{graph},a,b,\mathbf{B})$
  1. 1.
    Initialize an empty list reachable.
  2. 2.
    For each node $c\in\textit{graph.getAdjacentNodes}(b)$:
    1. (a)
      If $c=a$, continue.
    2. (b)
      If $\texttt{reachable}(\textit{graph},a,b,c,\mathbf{B})$ is true, append $c$ to reachable.
  3. 3.
    return reachable.
Helper: $\texttt{reachable}(\textit{graph},a,b,c,\mathbf{B})$
  1. 1.
    Let $\textit{collider}\leftarrow\textit{graph.isDefCollider}(a,b,c)$.
  2. 2.
    If $(\lnot\textit{collider}\;\lor\;\textit{graph.isUnderlineTriple}(a,b,c))$ and $b\notin\mathbf{B}$, return true.
  3. 3.
    Else return $\textit{collider}\wedge\textit{isAncestorOfAnyB}(b,\mathbf{B})$.
Algorithm 3 getReachableNodes and reachable
1:Goal: Build a candidate conditioning set $\mathbf{B}$ that blocks all blockable $m$-connecting paths from $x$ to $y$, ignoring the direct edge xyx\!-\!y on the first hop. If no valid blocking set exists, return Null.
2:Input: PAG $G$; nodes $x,y$; initial conditioning set containing (typically $\varnothing$).
3: Set $\mathbf{B}\leftarrow\textit{containing}$.
4: Initialize $\textbf{P}\leftarrow\langle x\rangle$.
5:for all nodes $b\in\mathrm{Adj}_{G}(x)$ do
6:  if $b=y$ then
7:   continue {Ignore the direct edge $x\text{--}y$ on the first hop}
8:  end if
9:  $r\leftarrow\texttt{find\_path\_to\_target}(G,\,a{=}x,\,b,\,y,\,\textbf{P},\,\mathbf{B})$
10:  if $r=\textsc{Unblockable}$ or $r=\textsc{Indeterminate}$ then
11:   return Null {No valid graphical sepset exists}
12:  end if
13:end forSTATE return $\mathbf{B}$ {All explored branches were blocked; may be $\varnothing$}
Algorithm 4 Recursive Blocking Procedure: block_paths_recursively
1
Input: Dataset $\mathcal{D}$ with $n$ variables
Significance level $\alpha$
Edge-minimal Markovian CPDAG procedure $\mathcal{M}$ for initial graph estimation
Conditional independence oracle
2
Output: Estimated PAG $\mathcal{G}$
3
4Step 1: Initial Graph Estimation
5 Obtain an initial graph $\mathcal{G}_{\text{CPDAG}}$ using edge-minimal Markovian CPDAG procedure $\mathcal{M}$
6 Initialize PAG $\mathcal{G}_{\text{PAG}}$ by copying $\mathcal{G}_{\text{CPDAG}}$
7
8Step 2: Extra Edge Removal Step
9 foreach edge XYX*\!\!{-}\!\!*Y in $\mathcal{G}_{\text{PAG}}$ do
10 foreach set $S$ where $S\subseteq\text{adj}(X,\mathcal{G}_{\text{PAG}})\setminus\{Y\}$ or $S\subseteq\text{adj}(Y,\mathcal{G}_{\text{PAG}})\setminus\{X\}$ do
11    if $X{\;\perp\!\!\!\perp\;}Y\mid S$ then
12         Remove edge XYX*\!\!{-}\!\!*Y from $\mathcal{G}_{\text{PAG}}$
13         Store $S$ in $\mathcal{S}(X,Y)$
14       break
15
16
17
18
19Step 3: Revised R0 Rule (Collider Identification)
20 Orient each edge XYX*\!\!{-}\!\!*Y in $\mathcal{G}_{\text{PAG}}$ as $X\circ-\circ Y$
21
22foreach unshielded triple $\langle X,Y,Z\rangle$ in $\mathcal{G}_{\text{PAG}}$ do
23 if $X\rightarrow Y\leftarrow Z$ in $\mathcal{G}_{\text{CPDAG}}$ then
24      Orient as $X*\to Y\leftarrow*Z$ in $\mathcal{G}_{\text{PAG}}$
25
26 else if $\sim adj(X,Y,\mathcal{G}_{\text{CPDAG}})$ then
27    $S\leftarrow\mathcal{S}(X,Y)$
28    if $S\neq\textsc{Null}$ and $Y\notin S$ then
29         Orient as $X*\to Y\leftarrow*Z$ in $\mathcal{G}_{\text{PAG}}$
30       break
31
32
33
34
35Step 4: Final Collider Orientation and FCI Rules
36 Apply FCI final orientation rules to refine $\mathcal{G}_{\text{PAG}}$
37
38return $\mathcal{G}_{\text{PAG}}$
Algorithm 5 *-FCI Algorithm (“Star-FCI”)
1Goal: Decide whether there exists an unblocked path from node $a$ to node $y$ passing through node $b$, given a dynamically maintained conditioning set $\mathbf{B}$.
2Return values: Null if no valid blocking set can be found; otherwise $\mathbf{B}$ (possibly $\varnothing$).
3Inputs:
  • graph: The graph structure
  • a,b,y: Node identifiers
  • P: Set of nodes already visited on this branch
  • $\mathbf{B}$: Current conditioning set (modifiable)
  • F: Optional set of vertices to exclude from traversal (e.g., fixed colliders on discriminating paths)
Procedure:
  1. 1.
    Target check. If $b=y$, return UNBLOCKABLE.
  2. 42.
    Cycle guard. If $b\in\textbf{P}$, return UNBLOCKABLE.
(A revisitation means we cannot certify blocking on this branch.)
  • 3.
    Add $b$ to the path: insert $b$ into P. 4. Compute neighbors to explore: $\textit{reachable}\leftarrow\texttt{getReachableNodes}(\textit{graph},a,b,\mathbf{B})\setminus\textbf{P}\setminus\textbf{F}.$
  • 5.
    Case 1: $b$ is latent or $b\in\mathbf{B}$.
    1. (a)
      For each $c\in\textit{reachable}$:
      • $\textit{res}\leftarrow\textit{find\_path\_to\_target}(\textit{graph},b,c,y,\textbf{P},\mathbf{B},\textbf{F})$.
      • If $\textit{res}=\texttt{UNBLOCKABLE}$: remove $b$ from P; return UNBLOCKABLE.
    2. (b)
      All continuations blocked: remove $b$ from P; return BLOCKED.
  • 56.
    Case 2: $b$ is neither latent nor in $\mathbf{B}$.
    1. (a)
      Branch A (do not add $b$ to $\mathbf{B}$).
      • For each $c\in\textit{reachable}$:
        • $\textit{res}\leftarrow\textit{find\_path\_to\_target}(\textit{graph},b,c,y,\textbf{P},\mathbf{B},\textbf{F})$.
        • If $\textit{res}=\texttt{UNBLOCKABLE}$: remove $b$ from P; return UNBLOCKABLE.
    2. (b)
      Branch B (temporarily add $b$ to $\mathbf{B}$).
      • Insert $b$ into $\mathbf{B}$.
      • Recompute
        $\textit{reachable}\leftarrow\texttt{getReachableNodes}(\textit{graph},a,b,\mathbf{B})\setminus\textbf{P}\setminus\textbf{F}.$
      • For each $c\in\textit{reachable}$:
        • $\textit{res}\leftarrow\textit{find\_path\_to\_target}(\textit{graph},b,c,y,\textbf{P},\mathbf{B},\textbf{F})$.
        • If $\textit{res}=\texttt{UNBLOCKABLE}$: remove $b$ from $\mathbf{B}$; remove $b$ from P; return UNBLOCKABLE.
      • Remove $b$ from $\mathbf{B}$ (backtrack).
    3. (c)
      Conclude Case 2: remove $b$ from P; return BLOCKED.
  • Return Value: UNBLOCKABLE if some $a\!\leadsto\!y$ path through $b$ cannot be blocked under the explored choices; BLOCKED if all such paths are blocked by $\mathbf{B}$ (with or without temporarily conditioning on $b$).
    Algorithm 6 find_path_to_target Procedure
    1
    Input : PAG $G=(V,E)$;
    distinct vertices $x,y\in V$;
    (optional) oracle $\mathcal{M}$ providing $m$-separation in $G$;
    independence test $\mathcal{T}$;
    orientation helper $\mathcal{F}$;
    maximum blocking-path length $L_{b}$;
    maximum discriminating-path length $L_{d}$;
    subset-depth bound $d$ (1-1 = no bound)
    Output : Separating set $S\subseteq V$: in oracle mode $x\mathrel{\perp\!\!\!\perp_{m}}_{G}y\mid S$; in test mode $x{\;\perp\!\!\!\perp\;}y\mid S$; or Null if none exists
    2
[…]
    4 $\mathcal{F}.\mathrm{finalOrientation}(G)$;
    5 $\mathcal{F}.\mathrm{setDoR4}(\mathrm{true})$;
    6
[…]
    8
[…]
    10
[…]
    12 $d_{1}\leftarrow\begin{cases}|NF_{\mathrm{cand}}|,&d=-1\\ d,&\text{otherwise}\end{cases}$
[…]
    14 $d_{2}\leftarrow\begin{cases}|\textit{Common}|,&d=-1\\ d,&\text{otherwise}\end{cases}$
    15foreach $NF\subseteq NF_{\mathrm{cand}}$ with $|NF|\leq d_{1}$ do
    16 $B\leftarrow\mathrm{block\_paths\_recursively}(G,x,y,\emptyset,NF,L_{b})$
    17
    18 foreach $C\subseteq\textit{Common}$ with $|C|\leq d_{2}$ do
    19    if $\exists c\in C:\;x\to c\leftarrow y$ then // skip definite colliders
    20       continue
    21
    22    $S\leftarrow B\setminus C$
    23
    24    if $\mathcal{M}\neq\varnothing$ then
           $\textit{indep}\leftarrow\mathcal{M}.\mathrm{markovIndependence}(x,y,S)$
           // $x\mathrel{\perp\!\!\!\perp_{m}}_{G}y\mid S$
    25
    26    else
           $\textit{indep}\leftarrow\mathcal{T}.\mathrm{checkIndependence}(x,y,S)$
           // $x{\;\perp\!\!\!\perp\;}y\mid S$
    27
    28
    29    if indep then
    30       return $S$
    31
    32
    33return Null
    Algorithm 7 R4 Discriminating-Path Optimization via Recursive Blocking
    A.2 Enumerating Discriminating and Pre-Discriminating Paths
    Setup and conventions.
    Let $G$ be a PAG (or partially oriented graph) over measured variables. We treat paths as ordered tuples $\langle v_{0},\dots,v_{k}\rangle$ (not bold), and we boldface node sets (e.g., $\mathbf{C},\mathbf{S}$). We assume access to the usual adjacency/orientation queries: $\operatorname{Adj}_{G}(\cdot)$, $\text{isParentOf}(\cdot,\cdot)$, $\text{getEndpoint}(\cdot,\cdot)\in\{\textsc{Tail},\textsc{Circle},\textsc{Arrow}\}$, and $\text{nodesInto}(t,\textsc{Arrow})=\{x:x*\!\!\rightarrow t\}$.
    Definition 2 (Pre-discriminating and discriminating paths).
    Fix distinct nodes $w,y$ with $w\!\rightarrow y$ (strict) or $w*\!\!\rightarrow y$ (relaxed). A pre-discriminating path for $\langle w,y\rangle$ is a path $\langle x,\dots,v,y\rangle$ whose internal nodes (between $v$ and $x$) are all colliders and parents of $y$, and whose penultimate edge satisfies $v\circ\!\!\rightarrow y$. A discriminating path additionally satisfies the usual FCI/PAG non-adjacency side condition between the endpoints of certain triples (as in the standard R4/R5 usage); we delegate the subtle case checks to a predicate $\textsc{ExistsDiscriminatingPath}(x,w,v,y,G)$.
    The routine below enumerates all such paths, with two modes: (i) strict (checkXYNonadjacency=true) requires $w\!\rightarrow y$; (ii) relaxed allows a covering edge $w*\!\!\rightarrow y$ but forbids $y\!\rightarrow w$. We also bound search length by maxLen (set 1-1 for no bound).
[…]
    2:for all $w\in V(G)$ do
    3:  for all $y\in\operatorname{Adj}_{G}(w)$ do
    4:   $\mathbf{Out}\leftarrow\mathbf{Out}\cup\ \textsc{ListDiscriminatingPaths}(G,w,y,\texttt{maxLen},\texttt{checkXYNonadjacency})$
    5:  end for
    6:end for
    7:return $\mathbf{Out}$
[…]
    2:if checkXYNonadjacency and $\neg(w\!\rightarrow y)$ then
    3:  return $\mathbf{Out}$
    4:end if
[…]
    6:  if $y\!\rightarrow w$ then
    7:   return $\mathbf{Out}$
    8:  end if
    9:end if
[…]
    11:for all $v\in\mathbf{V_{cand}}$ do
    12:  if $v\in\{w,y\}$ then
    13:   continue
    14:  end if
    15:  if $y\circ\!\!\rightarrow v$ and $v\!\rightarrow y$ then
    16:   $\textsc{DiscBFS}(G,w,v,y,\texttt{maxLen},\texttt{checkXYNonadjacency},\mathbf{Out})$
    17:  end if
    18:end for
    19:return $\mathbf{Out}$
[…]
    1: Initialize FIFO queue $Q$
    2: Enqueue state $(t{=}w,\;p{=}\textsc{Null},\;\textit{body}{=}\langle\,\rangle)$
    3:while $Q$ not empty do
    4:  Pop $(t,p,\textit{body})$
    5:  if $p\neq\textsc{Null}$ then
    6:   if $p\not\!\!\rightarrow t$ then
    7:    continue {collider at tt}
    8:   end if
    9:   if $\neg(t\!\rightarrow y)$ then
    10:    continue {interior node must parent yy}
    11:   end if
    12:  end if
    13:  for all $x\in\text{nodesInto}(t,\textsc{Arrow})$ do
    14:   if ($x=p\lor(x\in\textit{body}$) then
    15:    continue
    16:   end if
    17:   $\textit{body}^{\prime}\leftarrow\textit{body}$ appended with $t$
    18:   if ($\texttt{maxLen}\neq-1$) ($|\textit{body}^{\prime}|>\texttt{maxLen}$) then
    19:    continue
    20:   end if
    21:   Define candidate path $\pi\leftarrow\langle x,\;\textit{body}^{\prime},\;v,\;y\rangle$
    22:   if $\textsc{ExistsDiscriminatingPath}(x,w,v,y,\pi,G,\texttt{checkXYNonadjacency})$ then
    23:    $\mathbf{Out}\leftarrow\mathbf{Out}\cup\{\pi\}$
    24:   end if
    25:   if $x\!\rightarrow y$ then
    26:    Enqueue $(t{=}x,\;p{=}t,\;\textit{body}{=}\textit{body}^{\prime})$
    27:   end if
    28:  end for
    29:end while
[…]
    1:Input: candidate path $\pi=\langle x,\ldots,v,y\rangle$; PAG $G$; flag checkXYNonadjacency
    2:Output: true if $\pi$ forms a valid discriminating path in $G$
[…]
    4:  return false
    5:end if
    6:if checkXYNonadjacency and $x$ adjacent to $y$ then
    7:  return false
    8:end if
    9:if $v$ not adjacent to $y$ then
    10:  return false
    11:end if
    12:if $w\notin$ interior nodes of $\pi$ then
    13:  return false
    14:end if
    15: Let $\mathbf{p}\leftarrow$ reversed interior of $\pi$ with $x$ prepended and $v$ appended
    16:for $i=1$ to $|\mathbf{p}|{-}2$ do
    17:  $(n_{1},n_{2},n_{3})\leftarrow(\mathbf{p}[i{-}1],\mathbf{p}[i],\mathbf{p}[i{+}1])$
    18:  if not $\textsc{DefCollider}(n_{1},n_{2},n_{3},G)$ then
    19:   return false
    20:  end if
    21:  if checkXYNonadjacency and $n_{2}\not\!\!\rightarrow y$ then
    22:   return false
    23:  end if
[…]
    25:   return false
    26:  end if
    27:end for
    28:if $\neg v\rightarrow w$ then
    29:  errorvv must point to $w$
    30:end if
    31:return true
[…]
    Remarks and guarantees.
    • Strict vs. relaxed mode. When checkXYNonadjacency is true, we enforce $w\!\rightarrow y$; when false, we allow $w*\!\!\rightarrow y$ but reject $y\!\rightarrow w$. This matches the two modes in our implementation.
    • Interior-collider/parent invariant. Lines 6–8 in Alg. 10 enforce that every interior node on the upstream body is a collider and a parent of $y$, which is exactly the structure required by (pre-)discriminating paths used by FCI rules.
    • Termination and complexity. The BFS is acyclic on each branch (we forbid repeats in body) and can be bounded by maxLen. In the worst case the search is $O\!\left(\sum_{t}\deg^{+}(t)\right)$ per seed with pruning, and typically much faster due to the parent-of-$y$ filter.
    • Validation predicate. The routine $\textsc{ExistsDiscriminatingPath}(\cdot)$ encapsulates the FCI side-conditions (e.g., uncoveredness/non-adjacency of certain triples, endpoint patterns). Keeping these checks local isolates graph-specific details from the traversal.
    • Output. $\mathbf{Out}$ is a set of (pre-)discriminating paths $\pi$ (tuples). No node sets are returned here; all set-valued worklists or outputs are boldfaced.
    Appendix B Correctness of the Recursive Blocking Procedure
    B.1 Soundness (formal fix-point lemma)
    B.2 Soundness (formal fix-point lemma)
    Let $\mathtt{block\_paths\_recursively}(G,x,y,\textit{containing},\textit{notFollowed},L)$ denote the recursive blocking procedure with a maximum path-length limit $L$ (i.e., paths longer than $L$ edges are not explored).
    Theorem 5 (Soundness of Recursive Blocking).
    Let $G$ be a finite mixed ancestral graph (MAG). Let $\mathtt{block\_paths\_recursively}(G,x,y,\textit{containing},\textit{notFollowed},L)$ return a set $B$ (i.e., it terminates without returning Null). Then
    $\operatorname{Adj}(x,y,G)\ \ \vee\ \ \forall p\,\bigl[\operatorname{Path}(x,y,p,G)\ \Rightarrow\ \operatorname{Blocked}(p,B)\bigr],$
    where $\operatorname{Blocked}(p,B)$ is $m$-blocking by $B$ in $G$, and $\operatorname{Adj}(x,y,G)$ is adjacency in $G$.
    Proof sketch (via inducing paths).
    Assume the procedure halts and returns B, and suppose for contradiction that $\neg\operatorname{Adj}(x,y,G)$ and there exists an x–y path p that is open given B. We show that p must then be an inducing path, which in a MAG implies $\operatorname{Adj}(x,y,G)$—a contradiction.
    Step 1 (All interiors on p are colliders). Let $p=\langle x=v_{0},v_{1},\ldots,v_{k}=y\rangle$ be open given B. If some interior node $v_{i}$ were a noncollider on p with $v_{i}\notin B$, then the triple $(v_{i-1},v_{i},v_{i+1})$ would be traversable under the algorithm’s reachable predicate and would force the recursive call from the first hop $x\to v_{1}$ ultimately to return Unblockable unless $v_{i}$ were added to B. Since the final outcome is that the routine halts without ever returning Unblockable, this cannot happen. Hence every interior node of p is a collider or lies in B. But if an interior noncollider were in B, p could not be open given B. Therefore, every interior node on p is a collider.
    Step 2 (No activation via descendants; colliders not in B). Suppose a collider v on p were activated via a strict descendant $d\in B$. Along the directed segment $v\!\to\!\cdots\!\to d$, let u be the first noncollider encountered on the witnessed continuation. If the segment had length 1 ($v\!\to d$), then d would be a direct noncollider child of v; in that case the continuation would already be blocked once d entered B, and v would not be activated via d. Hence the segment must have length $\geq 2$, so such a u exists. By the algorithm’s minimality invariant—it adds each necessary noncollider as soon as it is encountered on an open continuation—u would have been added to B before d, contradicting minimality. Hence no collider on p is activated via a strict descendant in B. Moreover, by invariant, the procedure never adds colliders to B. Therefore colliders on p are not activated by membership in B.
    Step 3 (Inducing path). By Step 1, all interiors of p are colliders. By Step 2 and the invariant that B contains only noncolliders, each interior collider must be an ancestor of an endpoint (or of a selection node, if present). Hence every interior node is a collider in $\mathrm{An}(\{x,y\}\cup S)$, so p is an inducing path.
    Step 4 (Inducing path implies adjacency in a MAG). By maximality of MAGs, if there exists an inducing path between two distinct vertices, those vertices are adjacent. Thus the existence of the open path p (which we have shown must be inducing) implies $\operatorname{Adj}(x,y,G)$, contradicting our assumption $\neg\operatorname{Adj}(x,y,G)$.
    Therefore, if the routine halts with B and x and y are nonadjacent, no x–y path remains open given B. Equivalently, $\forall p\,(\operatorname{Path}(x,y,p,G)\Rightarrow\operatorname{Blocked}(p,B))$. ∎
    Corollary 3 (Adjacency Case).
    Under the assumptions of Theorem 1, suppose $\mathtt{block\_paths\_recursively}(G,x,y,\dots)$ halts normally (returning a non-null set B) and that x and y are adjacent in G. Then every non-inducing x–y path in G is m-blocked by B:
    $\operatorname{Adj}(x,y,G)\ \wedge\ B\neq\textsc{Null}\ \Longrightarrow\ \forall p\,\bigl[\operatorname{Path}(x,y,p,G)\ \wedge\ \neg\operatorname{Inducing}(p)\ \Rightarrow\ \operatorname{Blocked}(p,B)\bigr].$
    Proof sketch, When $\operatorname{Adj}(x,y,G)$, there exists at least one inducing path between x and y (the path defining the adjacency). We must show that every non-inducing x–y path p is m-blocked by the final blocking set B returned by the procedure.
    Step 1 (Exploration over neighbors of x). For each neighbor $b\in\operatorname{Adj}_{G}(x)$ with $b\neq y$, the algorithm invokes $\texttt{find\_path\_to\_target}(x,b,y,\dots)$ to explore every admissible continuation from x through b toward y, subject to the same reachable predicate used in the nonadjacent case. By the reachability completeness invariant, every potential open continuation of a non-inducing x–y path will eventually be examined.
    Step 2 (Blocking of noncollider continuations). Whenever the exploration encounters a triple (a,v,c) in which v is a noncollider not yet in B, the algorithm—by its noncollider-only addition rule—either (i) adds v to B to block that continuation, or (ii) returns Unblockable. Since the assumption here is that the routine halts normally with non-null B (i.e., no branch ever returned Unblockable), case (ii) never occurs. Therefore, every reachable noncollider continuation is eventually closed by adding the relevant noncollider to B.
    Step 3 (Halting and completeness). By the halting condition, termination implies that no further reachable triple with a noncollider $v\notin B$ remains open between x and y. Hence, after halting, every $x{\text{–}}y$ path whose interior nodes include at least one noncollider is m-blocked by the final B.
    Step 4 (Remaining open paths are inducing). The only x–y paths that can remain m-open given B are those whose interior nodes are all colliders and each such collider is an ancestor of an endpoint. These are precisely the inducing paths in G. In a MAG, inducing paths correspond exactly to adjacencies.
    Step 5 (Conclusion). Therefore, when $\operatorname{Adj}(x,y,G)$ and the routine halts with non-null B, every non-inducing x–y path is m-blocked by B, while any remaining open path must be inducing, accounting for the observed adjacency between x and y. ∎
    Remark 1 (Generality of the result).
    Although Theorem 5 and Corollary 3 are stated for MAGs, the recursive blocking argument applies unchanged to DAGs, CPDAGs, and PAGs. In DAGs and CPDAGs, $m$-separation reduces to $d$-separation, and in PAGs the same reasoning holds with mixed endpoints.
    B.3 Completeness (over the FCI search space)
    Lemma 2 (Domain restriction).
    Let $G$ be a PAG/MAG and fix $x\neq y$. Every vertex $b$ that $\mathtt{block\_paths\_recursively}(G,x,y,\dots)$ adds to $B$ lies in the standard FCI search space $\mathcal{S}(x,y)=\operatorname{Adj}(x)\cup\operatorname{Adj}(y)\cup\operatorname{PossibleDsep}(x,y)$.
    Proof sketch.
    Recursive blocking only proposes conditioning on vertices that occur as noncolliders on paths it actively traverses from $x$ toward $y$. Every interior node on such a traversed path is either a collider or an ancestor of a collider (otherwise the traversal would stop), and triangles are only crossed when the middle node is not a definite noncollider. This is precisely the inclusion criterion for $\operatorname{PossibleDsep}(x,y)$ in a PAG. Neighbors of $x$ or $y$ are in $\operatorname{Adj}(x)$ or $\operatorname{Adj}(y)$. Hence any $b$ considered by recursive blocking lies in $\mathcal{S}(x,y)$. ∎
    Lemma 3 (Termination).
    Fix $x\neq y$ and let $U=\mathcal{S}(x,y)$. Along any run of $\mathtt{block\_paths\_recursively}(G,x,y,\dots)$ with no path-length cap and $\textit{notFollowed}=\varnothing$, the sequence of conditioning sets $B_{0}\subset B_{1}\subset\cdots$ produced by the algorithm has strictly increasing $B_{i}$ and is bounded by $U$. Hence the run performs at most $|U|$ additions and halts.
    Proof sketch.
    By construction, RB adds only fresh vertices (never removes), so $B$ grows strictly when it grows. By Lemma 2, all additions lie in $U$, which is finite. Thus there is no infinite sequence of proper inclusions. ∎
    Corollary 4 (Completeness over $\mathcal{S}$).
    Let $U=\mathcal{S}(x,y)$ and assume the oracle setting with no path-length cap and $\textit{notFollowed}=\varnothing$. If there exists $S\subseteq U$ with $\text{m-sep}(x,y\mid S)$, then $\mathtt{block\_paths\_recursively}(G,x,y,\dots)$ halts and returns some $B\subseteq U$ with $\text{m-sep}(x,y\mid B)$.
    Proof sketch.
    By Lemma 3, it must halt after finitely many additions. At halt, Theorem 5 (soundness) shows that all xxyy paths are blocked by $B$, i.e. $\text{m-sep}(x,y\mid B)$. By Lemma 2, $B\subseteq U$. ∎
    B.4 Certification Completeness
    Proposition 1 (Certification completeness under Markovness).
    Assume the oracle setting where $G$ is a MAG or PAG Markov to the true distribution $\mathbb{P}$ (with no Faithfulness assumed), and that $\mathtt{block\_paths\_recursively}(G,x,y,\dots)$ halts with blocking set $B\subseteq\mathcal{S}(x,y)$. Let $C_{xy}=\operatorname{Adj}_{G}(x)\cap\operatorname{Adj}_{G}(y)$ denote the common adjacents of $x$ and $y$, and define
    $\mathcal{F}_{xy}=\{\,B\setminus D^{\prime}\mid D^{\prime}\subseteq C_{xy}\,\}.$
    Then, if there exists any conditioning set $S$ (not necessarily contained in $\mathcal{S}(x,y)$) such that $X{\;\perp\!\!\!\perp\;}Y\mid S$ in $\mathbb{P}$, there exists some $S^{\prime}\in\mathcal{F}_{xy}$ with $X{\;\perp\!\!\!\perp\;}Y\mid S^{\prime}$.
    Proof sketch.
    (i) RB baseline. By Theorem 5, the blocking set $B$ $m$-separates $x$ and $y$ in $G$. Each variable is added to $B$ only when required to block a reachable open noncollider continuation, ensuring that $B$ is locally minimal with respect to all explored $x\text{--}y$ paths.
    (ii) Essential vs. dispensable variables. Every $v\in B\setminus C_{xy}$ was added to block a specific non-inducing $m$-connecting path. Removing such a $v$ would reopen that path and destroy separation. Hence, any valid separating set for $(x,y)$ must include all of these “essential” vertices. By contrast, vertices in $C_{xy}$ may not lie on any active non-inducing path; they can sometimes be dropped without re-opening a connection.
    (iii) Completeness of $\mathcal{F}_{xy}$. The family $\mathcal{F}_{xy}$ enumerates all sets obtained by optionally omitting some or all of the potentially dispensable common neighbors $C_{xy}$ from $B$, while retaining all essential blockers. Thus, if there exists any true separating set $S$ in $\mathbb{P}$, it must coincide with or be a subset of some member of $\mathcal{F}_{xy}$, since any smaller set lacking an essential blocker would fail to $m$-separate $x$ and $y$.
    (iv) Markov-consistent certification. Because $G$ is Markov to $\mathbb{P}$, every $S^{\prime}\in\mathcal{F}_{xy}$ that $m$-separates $x$ and $y$ also satisfies $X{\;\perp\!\!\!\perp\;}Y\mid S^{\prime}$ in $\mathbb{P}$. Hence, if $X{\;\perp\!\!\!\perp\;}Y\mid S$ holds in the true distribution, there exists a corresponding $S^{\prime}\in\mathcal{F}_{xy}$ that certifies the same independence. Deleting the edge xyx*\!-\!*y or orienting the induced collider preserves Markovness and prepares the graph for the subsequent FCI orientation phase. ∎
    B.5 Effects of Unfaithfulness
    The preceding result assumes only Markovness, not Faithfulness. In the oracle setting this is sufficient for correctness: every conditional independence detected by recursive blocking corresponds to a genuine $m$-separation in the underlying MAG, and each certified separating set $S^{\prime}\in\mathcal{F}_{xy}$ therefore reflects a true structural independence. However, in finite-sample or unfaithful situations, apparent independences may arise that are not $m$-separations in any causal MAG.
    An unfaithful independence occurs when two variables are statistically independent only because of numerical parameter cancellation rather than structural separation. If FCIT acts on such a spurious independence—by deleting an edge or orienting a collider—the resulting graph may violate the ancestral or maximal properties that define a legal PAG. In this case no valid MAG can realize the asserted independence, and the resulting structure ceases to represent a Markov equivalence class.
    To ensure interpretability of search results, the implementation therefore includes a legality reversion step: whenever an update would yield an illegal PAG, the modification is reverted and the current graph is projected back into the space of legal PAGs. This step has no effect under Faithfulness, where all detected independences are structural and the procedure preserves legality automatically, but it provides robustness in the presence of unfaithful or noisy data by preventing the search from producing non-representable graphs.
    B.6 Relationship to Zhang’s m-Separation
    Algorithms 4 and 6 operate on a PAG or MAG $G=(V,E)$, two distinct vertices $x,y\in V$, a set of mandatory conditioning vertices $C\subseteq V$ (containing), and a set of vertices that may not be traversed $F\subseteq V$ (notFollowing). They return a blocking set $B$ obeying $C\subseteq B\subseteq V\setminus F$.
    A direct edge xyx*\!-\!*y has no interior vertices and therefore provides no opportunity to satisfy the blocking clauses by conditioning on a noncollider or leaving a collider unactivated. Hence, while clause (ii) (“all colliders are in or have descendants in BB”) holds trivially, this does not imply independence as long as $x$ and $y$ remain adjacent. To certify $m$-separation, both adjacency and path conditions must be satisfied. Zhang’s criterion [17] explicitly requires
    $\neg\operatorname{Adj}(x,y,G)\quad\text{and}\quad\forall p\,\operatorname{Blocked}(p,B),$
    which matches the structure of our Soundness theorem. Once the edge xyx*\!-\!*y is removed, Soundness supplies the second clause, ensuring that all $x\text{--}y$ paths are blocked. The following observations then establish constructive equivalence:
    1. (i)
      Once $x$ and $y$ become nonadjacent, the Fix-point Lemma ensures that the current blocking set $B$ blocks every $x\text{--}y$ path; hence $\text{m-sep}(x,y\mid B)$.
    2. (ii)
      Conversely, if there exists a set $B^{\star}$ with $\text{m-sep}(x,y\mid B^{\star})$, the algorithm eventually halts while $x$ and $y$ are still nonadjacent. Monotone growth gives $C\subseteq B\subseteq B^{\star}$, and Soundness yields $\text{m-sep}(x,y\mid B)$. Thus the procedure always returns a valid separating set whenever one exists.
    Hence, whenever $x$ and $y$ are $m$-separable, the recursive blocking procedure finds a separating set consistent with the user’s constraints.
    B.7 Complexity
    Let $\Delta_{x}=|\mathrm{Adj}_{G}(x)|$ and let $\ell$ be the path-length cap used by the routine (maxPathLength). In each call, the recursion explores a binary branch (first without, then with $b\in\mathbf{B}$), and along any single branch the path guard prevents revisiting vertices. The cost of one branch is linear in the size of the explored subgraph (dominated by adjacency scans), i.e., $\mathcal{O}(|E|)$ in the worst case.
    Bounded depth.
    If $\ell$ is a fixed constant, the total number of branches is $\mathcal{O}(2^{\ell})$, so the overall complexity is
    $\mathcal{O}\!\left(\Delta_{x}\cdot 2^{\ell}\cdot|E|\right).$
    Since $2^{\ell}$ is constant, this is effectively $\mathcal{O}(\Delta_{x}\,|E|)$.
    Unbounded depth.
    If $\ell$ grows with the graph (no path-length cap), the worst-case number of branches can be exponential in the path length, yielding
    $\mathcal{O}\!\left(\Delta_{x}\cdot 2^{\Theta(\ell)}\cdot|E|\right),$
    which in the worst case is exponential in $|V|$.
    Space.
    The algorithm stores $\mathbf{B}$ and the current path, each of size $\mathcal{O}(|V|)$, plus the descendants index (precomputed once), which is $\mathcal{O}(|V|+|E|)$.
    B.8 Concrete Example: Peter’s MAG
    Consider the MAG in Figure 15 with the following edges:
    $x\;\to\;b,\quad b\;\to\;z,\quad z\;\leftrightarrow\;y,\quad x\;\leftrightarrow\;a,\quad a\;\leftrightarrow\;y,\quad a\;\to\;b.$
    We want to determine whether $x$ and $y$ are m-separable.
    1. Initial Conditioning Set.
    We begin with $B=\varnothing$. The recursive blocking driver repeatedly calls $\textit{find\_path\_to\_target}(x,y,B)$ to detect any m-unblocked path from $x$ to $y$.
    2. Checking Paths.
    A candidate path is
    $x\;\to\;b\;\to\;z\;\leftrightarrow\;y.$
    Since $b\to z$ and $z\leftarrow y$ (by virtue of $z\leftrightarrow y$), the node $z$ is a collider. A collider path is BlockedAfter by default if neither the collider nor its descendants are in $B$. Because $B=\varnothing$, the path is already inactive. Hence, the procedure has no need to add $b$. Likewise, the path $x\leftrightarrow a\leftrightarrow y$ or any variant is also BlockedAfter under $\varnothing$.
    3. No Node Added.
    The algorithm finds no m-unblocked path and therefore terminates without modifying $B$. Consequently, $B$ remains empty. Indeed, $\varnothing$ suffices to m-separate $x$ and $y$.
    4. Conclusion.
    This example shows that, with correct collider detection, the path $x\to b\to z\leftrightarrow y$ is recognized as BlockedAfter under $\varnothing$. A naive method that tries to block $z$ because of the noncollider path $x\rightarrow b\rightarrow z$ would incorrectly add $b$. By contrast, the proper recursive blocking procedure sees no reason to alter $B$, returning $\varnothing$. This illustrates the crucial role of the find_path_to_target method.
    Figure 15: Peter’s MAG example. Since $z$ is a collider on the path $x\to b\to z\leftrightarrow y$, it is already BlockedAfter under $\varnothing$. No additional conditioning is needed.
    Appendix C Optimizing the Zhang Rules
    In FCIT, the final FCI orientation rules are applied repeatedly, and some of these rules can become bottlenecks in practice. The Zhang rules are given in Section C.1. In particular, we focus on optimizing four rules, whose standard implementations are typically the most time-consuming:
    • Rule R4, which requires searching for discriminating paths and can involve conditioning on many subsets of adjacent nodes of varaibles.
    • Rules R5 and R9, both of which involve identifying all-circle-endpoint paths of potentially unbounded length.
    • Rule R10, which requires finding certain semidirected paths to finalize circle endpoints.
    Below, we give our optimizations for each of these.
    C.1 Zhang’s Final FCI Orientation Rules
    For completeness, we recap here the orientation rules (R0–R10) that Zhang [17] proves are both sound and complete for FCI in the presence of latent variables and selection bias. In these rules, $\alpha,\beta,\gamma,\theta$ denote generic vertices, and the symbol $\ast$” is a wildcard that can represent an arrowhead ($\rightarrow$), a tail (-), or a circle ($\circ$). Any rule that does not apply in a given configuration can be ignored.
    R0 (Unshielded collider):
    For every unshielded triple $\langle\alpha,\gamma,\beta\rangle$ (i.e. $\alpha$ and $\beta$ both adjacent to $\gamma$ but not to each other):
    • If $\gamma\not\in\text{Sepset}(\alpha,\beta)$, orient $\alpha\ast\to\gamma\leftarrow\!\ast\beta$.
    R1:
    If $\alpha\ast\to\beta\;\circ-\!\ast\gamma$ and $\alpha$ and $\gamma$ are not adjacent, orient $\beta\to\gamma$.
    R2:
    If either $\alpha\to\beta\ast\to\gamma$ or $\alpha\ast\to\beta\to\gamma$, and there is an unoriented edge $\alpha\ast-\!\circ\gamma$, orient $\alpha\ast\to\gamma$.
    R3:
    If $\alpha\ast\to\beta\leftarrow\!\ast\gamma$, and $\alpha\ast-\!\circ\theta\circ-\!\ast\gamma$, with $\alpha$ and $\gamma$ not adjacent, and $\theta\ast-\!\circ\beta$, then orient $\theta\ast\to\beta$.
    R4 (Discriminating path):
    Let $u=\langle\theta,\dots,\alpha,\beta,\gamma\rangle$ be a discriminating path between $\theta$ and $\gamma$ for $\beta$, and suppose $\beta\circ-\!\ast\gamma$ is currently unoriented. Then:
    • If $\beta\in\text{Sepset}(\theta,\gamma)$, orient $\beta\circ-\!\ast\gamma$ as $\beta\to\gamma$.
    • Otherwise orient $\langle\alpha,\beta,\gamma\rangle$ as $\alpha\leftrightarrow\beta\leftrightarrow\gamma$.
    R5:
    If $\alpha\circ-\!\circ\beta$ and there is an uncovered circle-only path $\langle\alpha,\ldots,\beta\rangle$ such that $\alpha$ and $\beta$ are not adjacent to one another’s immediate neighbors on that path, orient $\alpha\circ-\!\circ\beta$ (and every edge on the path) as undirected, i.e. $\alpha-\beta$.
    R6:
    If $\alpha-\beta\;\circ-\!\ast\gamma$, then orient $\beta\circ-\!\ast\gamma$ as $\beta-\gamma$.
    R7:
    If $\alpha-\beta\;\circ-\!\circ\gamma$ and $\alpha$ and $\gamma$ are not adjacent, orient $\beta\circ-\!\circ\gamma$ as $\beta-\gamma$.
    R8:
    If $\alpha\to\beta\to\gamma$ or $\alpha-\beta\to\gamma$, and $\alpha\circ\!\to\gamma$ is present, orient $\alpha\circ\!\to\gamma$ as $\alpha\to\gamma$.
    R9:
    If $\alpha\circ\!\to\gamma$ and there is an uncovered potentially-directed path $\langle\alpha,\beta,\theta,\ldots,\gamma\rangle$ such that $\beta$ and $\gamma$ are not adjacent, then orient $\alpha\circ\!\to\gamma$ as $\alpha\to\gamma$.
    R10:
    Suppose $\alpha\circ\!\to\gamma$, $\beta\to\gamma\leftarrow\theta$, and there exist uncovered potentially-directed paths $p_{1}:\alpha\to^{\ast}\beta$ and $p_{2}:\alpha\to^{\ast}\theta$, where the two neighbors of $\alpha$ on $p_{1}$ and $p_{2}$ (denote them $\nu$ and $\omega$) are distinct and not adjacent. In that case, orient $\alpha\circ\!\to\gamma$ as $\alpha\to\gamma$.
    C.2 Optimizing R4 Using Recursive Blocking
    A major bottleneck among the final orientation rules is Rule R4, which seeks discriminating paths of the form
    $X*\!\rightarrow\dots\leftrightarrow\;W\;\leftarrow\!\circ\;V^{\prime}\;\circ\!\rightarrow\;Y,$
    where each internal collider along the path is a parent of $Y$. In the standard approach [14], applying R4 often requires exploring many potential conditioning sets, which is time-consuming and can lead to inaccuracies in dense graphs.
    To mitigate this, we leverage the recursive blocking procedure (Section 4) to focus only on a small, targeted collection of conditioning sets. Specifically:
    1. 1.
      Enumerate discriminating paths. Identify all valid and unoriented discriminating paths from $X$ to $Y$. Because FCIT prunes many extraneous edges beforehand, the number of such paths is usually modest.
    2. 2.
      Collect collider candidates. For each path, record the possible collider nodes (including $V^{\prime}$ if its endpoint with $Y$ is a circle). These nodes form a small set of “candidate colliders” that we either include or exclude in our conditioning sets.
    3. 3.
      Apply recursive blocking. For each subset $Q$ of these candidate colliders, call the block_paths_recursively subroutine to build a minimal set $S$ that blocks all active paths from $X$ to YY—assuming each node in $Q$ is treated as a collider.
    4. 4.
      Test subsets of common adjacents. Let $C$ be the set of common adjacents of $X$ and $Y$. For each subset $C^{\prime}\subseteq C$, we form $S^{\prime}=S\setminus C^{\prime}$ (or vice versa, depending on how $Q$ is handled) and test whether $X{\;\perp\!\!\!\perp\;}Y\mid S^{\prime}$. If $X{\;\perp\!\!\!\perp\;}Y\mid S^{\prime}$ and $V^{\prime}\notin S^{\prime}$, we orient the relevant structure as a new collider (e.g., $W\leftrightarrow V^{\prime}\leftarrow Y$); otherwise, we orient $\alpha\circ\!\to\gamma$ as a noncollider.
    By confining the conditioning sets to those implied by the recursive blocking procedure on legitimate discriminating paths, we drastically reduce both the path search space and the number of conditional independence tests performed. We give pseudocode for this procedure in Algorithm 7.
    C.3 Shortest-Path Searches for R5 and R9
    Rules R5 and R9 require checking whether a path composed entirely of circle endpoints exists between two nodes. Naively, one might enumerate all such paths, but this is prohibitively expensive for larger graphs. Instead, we construct an auxiliary circle-only graph:
    • The node set is the same as in the original PAG.
    • We add an undirected edge XYX-Y in the auxiliary graph if and only if $X\circ\!\!-\!\!\circ Y$ in the PAG.
    We then run a shortest-path algorithm (e.g., BFS or Dijkstra’s, typically BFS with uniform edge weights) from $X$. If $Y$ is reachable in this auxiliary graph, then at least one all-circle path exists in the original PAG. Because most edges with definite tails or arrowheads are excluded, the auxiliary graph is often much sparser, allowing us to detect or rule out circle-only paths much more efficiently than exhaustive enumeration.
    C.4 Uncovered Potentially-Directed Path Checking for R10
    Rule R10 is among the costliest FCI orientation rules because it requires verifying the existence of uncovered, potentially-directed (PD) paths. The rule applies when $\alpha\circ\!\to\gamma,\quad$ $\beta\to\gamma\leftarrow\theta,$ and there exist uncovered PD paths $p_{1}:\alpha\to^{*}\beta$ and $p_{2}:\alpha\to^{*}\theta$ whose first hops from $\alpha$ are through distinct, nonadjacent neighbors $\mu$ and $\omega$. If so, the circle on $\alpha\circ\!\to\gamma$ is replaced by a tail ($\alpha\to\gamma$).
    To guarantee correctness, we use an explicit depth-first search that enforces both constraints: (i) every step must follow an edge with no arrowhead into the previous node (potentially-directed), and (ii) each triple must be unshielded (uncovered). Subproblems are memoized to avoid redundant traversals. This yields an exact implementation of R10 that ensures all orientations remain valid under the oracle semantics of Zhang’s rules.
    Appendix D Supplementary Traces
    Supplementary Trace 1: Recursive Blocking Example
    This section provides a worked example tracing the behavior of block_paths_recursively on a case where the score-based DAG includes a spurious edge that FCIT correctly removes through recursive blocking. This example shows the recursive blocking procedure removing an extraneous edge due to latent confounding.
    True model.
    The true causal DAG includes an unobserved confounder $L$ that influences both $Y$ and $Z$, and is given by:
    $X\rightarrow Y\leftarrow L\rightarrow Z\leftarrow W$
    Here, $L$ is latent. The observed variables are $X$, $Y$, $Z$, and $W$. The induced PAG over these observed variables has the following structure:
    • $X\rightarrow Y$
    • $W\rightarrow Z$
    • $Y\leftrightarrow Z$ (due to unmeasured confounding via $L$)
    XXYYZZWWLL
    Figure 16: True causal DAG including latent variable $L$ (dashed arrows). Only $X$, $Y$, $Z$, and $W$ are observed.
    Score-derived DAG.
    Since the score-based method cannot represent latent confounding directly, it introduces the edge $X\rightarrow Z$ to cover the collider X>Y<ZX*->Y<-*Z so that the graph may remain Markov to the data. The resulting DAG is:
    $X\rightarrow Y\rightarrow Z\leftarrow W,\quad X\rightarrow Z$
    XXYYZZWW
    Figure 17: DAG returned by score-based structure search. The edge $X\rightarrow Z$ does not belong in the true PAG and will be tested for removal.
    Blocking analysis.
    We apply block_paths_recursively($X$, $Z$) to test whether the edge $X\rightarrow Z$ can be removed.
    P Type Status under $B=\{Y\}$
    $X\rightarrow Z$ Direct edge Active
    $X\rightarrow Y\rightarrow Z$ Non-collider chain Blocked (since $Y\in B$)
    $X\rightarrow Z\leftarrow W$ Collider at $Z$ Blocked (since $Z\notin B$)
    Table 3: Paths from $X$ to $Z$ considered by block_paths_recursively. All non-collider paths are BlockedAfter by $B=\{Y\}$.
    The recursive blocking procedure returns:
    $\boxed{B=\{Y\}}$
    Next, FCIT computes the set of common adjacents $C=\text{Adj}(X)\cap\text{Adj}(Z)=\{Y\}$, and tests conditional independence over all subsets of $B\setminus D$ for $D\subseteq C$:
    Subset $D\subseteq C=\{Y\}$ Tested set $B\setminus D$ Independence result
    $\emptyset$ $\{Y\}$ Dependent (fails)
    $\{Y\}$ $\emptyset$ Independent (succeeds)
    Table 4: Conditional independence tests performed by FCIT for $X{\;\perp\!\!\!\perp\;}Z$. The independence is found given the empty set.
    Conclusion.
    Since $X{\;\perp\!\!\!\perp\;}Z\mid\emptyset$, FCIT removes the edge $X\rightarrow Z$. Because the separating set does not contain $Y$, the unshielded triple XYZX-Y-Z is oriented as a collider:
    $X\rightarrow Y\leftarrow Z$
    This demonstrates how FCIT successfully recovers latent-induced collider structure by combining recursive blocking with selective independence testing.
    Supplementary Trace 2: Discriminating Path Required for Edge Removal
    This example illustrates how a discriminating path must be used during the edge removal phase so that FCIT can correctly orient a collider and remove a spurious edge. Without including this, incorrect adjacencies and orientations persist in the output PAG.
    True DAG.
    The true DAG is in Figure 18 (a). Measured nodes are shown in rectangular boxes, and latent variables are shown in elliptical boxes.
    (a) True DAG
    (b) Induced PAG
    (c) FCIT without discriminating path edge removal step
    Figure 18: Graph visualizations generated using Tetrad. Left: the true DAG (including latent variables). Middle: the correct PAG over observed variables. Right: PAG returned by FCIT when the discriminating path edge removal rule is disabled.
    Key structure.
    In the induced PAG, the triple DD*LLM*M is initially unoriented. FCIT uses the discriminating path:
    $D\leftrightarrow H\leftrightarrow L\rightarrow M$
    to determine that $L$ is a collider in this triple. Orienting $D\rightarrow L\leftarrow M$ enables the recursive blocking procedure to find a separating set between $D$ and $M$, allowing the spurious edge to be removed.
    Table 5 shows a summary of the differences between the PAG found with the discriminating path edge removal step and the PAG found without it. Without that step, there are several orientation mistakes and one adjacency mistake.
    Edge Correct PAG Without Discriminating Path Step
    D – H $D\leftrightarrow H$ $H\to D$
    H – L $H\leftrightarrow L$ $H\to L$
    D – L $\to L DD$ o–o $L$
[…]
    S – L (absent) $S\to L$
    Other edges identical identical
    Table 5: Comparison of key edges in the PAG with and without using the discriminating path orientation step in edge removal. Differences arise due to missed collider orientations affecting recursive blocking decisions.
    Conclusion.
    This example demonstrates that discriminating path orientation must occur during edge removal for recursive blocking to correctly identify separating sets and return a valid PAG. This requirement is specific to FCIT. In FCI and GFCI, adjacencies $X!\text{–}!Y$ are removed prior to orientation, using subsets of adjacents or Possible-D-SEP sets. FCIT, by contrast, identifies separating sets by tracing paths recursively—an approach that depends on correct collider orientation during edge removal, including colliders identified via discriminating paths. This design enables a more efficient and accurate latent variable search: it leverages richer graphical structure during edge removal and reduces the number of conditional independence tests to those involving subsets of the common adjacents of $X$ and $Y$ after the initial CPDAG search, a much smaller search space.
    Take-Home Message.
    Together, these traces illustrate how FCIT integrates recursive blocking with discriminating path orientation during edge removal, yielding separating sets that are both valid and efficient to compute. By using more of the graph’s structure earlier in the process, FCIT achieves correctness guarantees while sharply reducing the search space for conditional independence tests.
    Appendix E Auxiliary Figures
    Appendix: 20-Node Results for All Average Degrees
    For completeness, we provide all figures for the 20-node case with average degrees 2 and 6 (Figs. 1931 and Figs. 3244, respectively). These exhibit trends consistent with the average-degree 4 case discussed in the main text. At lower densities (average degree 2) all algorithms recover sparser, more stable structures, whereas at higher densities (average degree 6) recall improves but orientation accuracy declines slightly.
    Figure 19: Adjacency Precision (AP) for 20-node graphs with average degree 2. Precision remains uniformly high across all algorithms (>0.93>0.93). BOSS-FCI GRaSP-FCI, and GFCI achieve near-perfect accuracy ($\approx 1.0$) throughout, with LV-Lite lagging behimd a bit. FCI is also slightly behind but still quite high. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 20: Adjacency Recall (AR) for 20-node graphs with average degree 2. All algorithms improve sharply with sample size, essentially in parallel. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 21: Arrowhead Precision (AHP) for 20-node graphs with average degree 2. LV-Lite lags at small samples, and FCI is behind overall; the other algorithms do quite well. FCI remains slightly lower but converges steadily. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 22: Arrowhead Recall (AHR) for 20-node graphs with average degree 2. Recall improves monotonically with NN for all algorithms, exceeding 0.7 at the largest sample size. FCI attains the highest overall recall, followed closely by FCIT, while LV-Lite and the GFCI-family algorithms show similar convergence. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 23: Arrowhead Precision for Common Adjacencies (AHPC) for 20-node graphs with average degree 2. Most algorithms fare quite well here, with FCI lagging behind. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 24: Arrowhead Recall for Common Adjacencies (AHRC) for 20-node graphs with average degree 2. FCI achieves the highest arrowhead recall (>0.9>0.9) at large $N$, with other algorithms somewhat lower. Out of the algorithms with good AHPC, FCIT performs the best. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 25: Tail Precision (TP) for 20-node graphs with average degree 2. Most algorithms are in a top group, with LV-Lite and FCI lagging behind. FCI trails slightly at larger sample sizes. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 26: Tail Recall (TR) for 20-node graphs with average degree 2. The algorithm all consistently improve with sample size, with FCIT having the largest tail recalls uniformly. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 27: Arrow Path Precision for 20-node graphs with average degree 2. This precision is excellent here for all algorithms except FCI as NN increases. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 28: Tail Path Precision for 20-node graphs with average degree 2. Precision varies strongly with NN; for large samples. LV-Lite is systematically behind. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 29: Bidirected Latent Path Precision for 20-node graphs with average degree 2. There is considerable variation here, though FCI is lower overall. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 30: Proportion of well-formed PAGs for 20-node graphs with average degree 2. There is good PAG precision for this regime, with LV-Lite and FCIT returning perfect scores, as predicted. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 31: Runtime (CPU milliseconds) for 20-node graphs with average degree 2. All algorithms run quickly in this regime. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 32: Adjacency Precision (AP) for 20-node graphs with average degree 6. All algorithms maintain high precision (>0.93>0.93) across sample sizes. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 33: Adjacency Recall (AR) for 20-node graphs with average degree 6. LV-Lite and FCIT show the strongest gains with sample size, reaching $\approx 0.7$ at large NN. GFCI, BOSS-FCI, and FCI improve more modestly, and GRaSP-FCI remains lowest. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 34: Arrowhead Precision (AHP) for 20-node graphs with average degree 6. LV-Lite attains the highest precision and improves with NN; FCIT is next best. GFCI and BOSS-FCI remain around $0.7$, GRaSP-FCI is lower, and FCI lowest overall. There is a consistent dip at N=500N{=}500. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 35: Arrowhead Recall (AHR) for 20-node graphs with average degree 6. All curves rise with NN; LV-Lite and FCIT are highest at large NN ($\approx 0.6$), while the GFCI-family methods and FCI trail. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 36: Arrowhead Precision for Common Adjacencies (AHPC) for 20-node graphs with average degree 6. LV-Lite maintains the highest precision and improves with NN; FCIT also increases steadily. BOSS-FCI and GFCI remain around $0.72$, GRaSP-FCI is lower, and FCI lowest overall. There is a consistent dip at N=500N{=}500. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 37: Arrowhead Recall for Common Adjacencies (AHRC) for 20-node graphs with average degree 6. Recall increases with NN for all methods. At large $N$, BOSS-FCI is highest (near $0.93$), followed by GFCI and FCI; FCIT and GRaSP-FCI are close behind, and LV-Lite is slightly lower. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 38: Tail Precision (TP) for 20-node graphs with average degree 6. GRaSP-FCI and BOSS-FCI, and FCIT outperform the other algorithms here, though none of the algorithms performs especially well. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 39: Tail Recall (TR) for 20-node graphs with average degree 6. At the all sample sizes LV-Lite and FCIT have the highest scores, while the others falls short. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 40: Arrow Path Precision for 20-node graphs with average degree 6. LV-Lite and FCIT achieve the highest precision, increasing to $\approx 0.96$ at large NN. BOSS-FCI and GFCI cluster around $0.80$; GRaSP-FCI is lower and FCI lowest overall. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 41: Tail Path Precision for 20-node graphs with average degree 6. There is considerable variation for all methods across sample sizes. GRaSP-FCI and BOSS-FCI outperform the others consistently. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 42: Bidirected Latent Path Precision for 20-node graphs with average degree 6. There is weak performance on this statistic here. Precision peaks near intermediate NN and eases slightly at the largest NN. FCIT attains the highest precision at mid-range, with other algorithms trailing. LV-Lite orients no bidirected edges, yielding undefined precision, so it is not plotted. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 43: Proportion of well-formed PAGs for 20-node graphs with average degree 6. LV-Lite and FCIT consistently produce valid PAGs (=1.0=1.0), confirming structural correctness. Other algorithms fall off, sometimes sharply. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 44: Runtime (CPU milliseconds) for 20-node graphs with average degree 6. GRaSP-FCI incurs the highest cost, growing steeply at large NN ($\sim\!600$ ms). FCIT is moderate, while LV-Lite, LV-Lite, FCI, and GFCI remain consistently faster. Results are averaged over graphs with 0, 4, and 8 latent common causes.

### v3

@@
 FCIT guarantees well-formed PAGs and achieves higher precision and efficiency across sample sizes.
-Finally, we propose a lightweight heuristic, LV-Lite (Latent Variable “Light”), which returns the PAG of the BOSS DAG (Directed Acyclic Graph).
-Though not strictly sound for latent confounding, LV-Lite often matches FCIT’s accuracy while running substantially faster.
-Simulations and real-data analyses show that BOSS-FCI and GRaSP-FCI provide robust baselines, FCIT yields the best balance of precision and reliability, and LV-Lite offers a fast, near-equivalent alternative.
+Finally, we propose a lightweight heuristic, LV-Dumb (Latent Variable “Dumb”), which returns the PAG of the BOSS DAG (Directed Acyclic Graph).
+Though not strictly sound for latent confounding, LV-Dumb often matches FCIT’s accuracy while running substantially faster.
+Simulations and real-data analyses show that BOSS-FCI and GRaSP-FCI provide robust baselines, FCIT yields the best balance of precision and reliability, and LV-Dumb offers a fast, near-equivalent alternative.
 Together, these methods demonstrate that targeted and score-guided strategies can dramatically improve the efficiency and correctness of latent-variable causal discovery.
@@
 In many real-world settings, latent variables or selection bias complicate this task, undermining the reliability of traditional discovery methods.
-The Fast Causal Inference (FCI) algorithm [14] provides a principled framework for such cases by applying conditional independence (CI) tests to subsets of adjacent variables (or Possible-D-SEP sets).
+The Fast Causal Inference (FCI) algorithm [spirtes2001causation] provides a principled framework for such cases by applying conditional independence (CI) tests to subsets of adjacent variables (or Possible-D-SEP sets).
 While correct in theory, this strategy is fragile in practice: testing many subsets amplifies stochastic variation in $p$-values, leading to spurious independence claims, missing or extra edges, and unstable orientations.
@@
 As $|\mathcal{S}|$ grows, the chance that some $S\in\mathcal{S}$ spuriously satisfies $\mathcal{T}(X{\;\perp\!\!\!\perp\;}Y\mid S)$ increases, inflating Type I errors and producing overly sparse graphs and erroneous orientations.
-Beyond sampling variability, another challenge is almost unfaithfulness [18, 15].
+Beyond sampling variability, another challenge is almost unfaithfulness [zhang2008detection, spirtes2014uniformly].
 Even when the Causal Markov and Faithfulness assumptions hold exactly, small parameter perturbations can produce distributions that are nearly unfaithful—where partial correlations or other dependence measures are close to zero but not exactly zero.
 In finite samples, such “almost unfaithful” parameterizations lead to unstable independence decisions: small statistical fluctuations can flip dependence judgments, causing orientation errors or spurious adjacencies.
-As shown by Zhang and Spirtes, [18], this phenomenon can make reliable causal inference difficult even for consistent algorithms, motivating procedures that are robust to weak dependencies or employ stronger assumptions such as $k$-triangle faithfulness [15].
+As shown by zhang2008detection, this phenomenon can make reliable causal inference difficult even for consistent algorithms, motivating procedures that are robust to weak dependencies or employ stronger assumptions such as $k$-triangle faithfulness [spirtes2014uniformly].
 We therefore treat almost unfaithfulness as a complementary concern to the repeated-testing problem—structural rather than statistical in origin, but equally capable of producing unstable PAGs in practice.
 Several extensions have sought to mitigate such issues.
-Order-independent FCI variants [5] stabilize orientations, and score-based methods [^1] Scores in Tetrad currently cover linear Gaussian and linear non-Gaussian models (using BIC-type criteria), multinomial models, mixed multinomial/linear Gaussian models, and, more recently, mixed multinomial/nonlinear post-nonlinear models via truncated basis-function scores [11].
-extend causal search to latent-variable models through likelihood-based criteria [16, 4].
-Hybrid approaches such as Greedy FCI (GFCI; [10]) combine score-based search with CI testing, but still rely on adjacency-based phases—i.e., during the adjacency-removal stage, testing all subsets of nodes adjacent to a pair—and thus remain vulnerable to repeated testing and almost unfaithfulness.
+Order-independent FCI variants [colombo2014order] stabilize orientations, and score-based methods [^1] Scores in Tetrad currently cover linear Gaussian and linear non-Gaussian models (using BIC-type criteria), multinomial models, mixed multinomial/linear Gaussian models, and, more recently, mixed multinomial/nonlinear post-nonlinear models via truncated basis-function scores [ramsey2025scalablecausaldiscoveryrecursive].
+extend causal search to latent-variable models through likelihood-based criteria [triantafillou2016score, bhattacharya2021differentiable].
+Hybrid approaches such as Greedy FCI (GFCI; [ogarrio2016hybrid]) combine score-based search with CI testing, but still rely on adjacency-based phases—i.e., during the adjacency-removal stage, testing all subsets of nodes adjacent to a pair—and thus remain vulnerable to repeated testing and almost unfaithfulness.
 Constraint-based methods, by contrast, have the advantage of being inherently nonparametric and making no distributional assumptions except for those of the selected conditional independence test, which motivates the option of keeping such testing in the algorithm.
 Recent advances in score-based search motivate new hybrids that combine the distributional flexibility of constraint-based methods with the efficiency of likelihood-based optimization.
-Best Order Score Search (BOSS; [2]) provides a scalable, permutation-based alternative to FGES, while GRaSP [9] offers a related variant.
+Best Order Score Search (BOSS; [andrews2019learning]) provides a scalable, permutation-based alternative to FGES, while GRaSP [lam2022greedy] offers a related variant.
 Substituting these into the GFCI framework yields BOSS-FCI and GRaSP-FCI—hybrid algorithms that preserve the nonparametric generality of constraint-based search while leveraging score-based initialization to improve stability and scalability.
-The original GFCI algorithm [10] is sound and complete in the population limit under the Causal Markov and Faithfulness assumptions, returning the correct Partial Ancestral Graph when provided with perfect independence information.
+The original GFCI algorithm [ogarrio2016hybrid] is sound and complete in the population limit under the Causal Markov and Faithfulness assumptions, returning the correct Partial Ancestral Graph when provided with perfect independence information.
 However, in finite samples, GFCI and its variants still depend on adjacency-based conditioning and thus remain susceptible to repeated testing and almost unfaithfulness.
@@
 Together with a well-formedness check that enforces the rules of a Partial Ancestral Graph (PAG)—a mixed graph representing the Markov equivalence class of maximal ancestral graphs consistent with possible latent variables and selection bias—this targeted strategy guarantees that every intermediate graph remains structurally consistent, meaning it obeys the ancestral, acyclicity, and maximality constraints required of a valid PAG.
-The result is greater efficiency, more stable orientations, and structural correctness not ensured by prior FCI variants [17, 8].
-Finally, we introduce a deliberately simple heuristic, LV-Lite (Latent Variables “Light”), which outputs the PAG implied by the BOSS DAG without performing any latent-specific conditional independence testing.
-Though it cannot orient bidirected edges, LV-Lite scales exceptionally well and, in extensive simulation studies, often matches or exceeds the empirical accuracy of more complex algorithms when compared to the true PAG.
-Together, these four algorithms—BOSS-FCI, GRaSP-FCI, FCIT, and LV-Lite—form a continuum.
+The result is greater efficiency, more stable orientations, and structural correctness not ensured by prior FCI variants [zhang2008completeness, hu2024fast].
+Finally, we introduce a deliberately simple heuristic, LV-Dumb (Latent Variables “Dumb”), which outputs the PAG implied by the BOSS DAG without performing any latent-specific conditional independence testing.
+Though it cannot orient bidirected edges, LV-Dumb scales exceptionally well and, in extensive simulation studies, often matches or exceeds the empirical accuracy of more complex algorithms when compared to the true PAG.
+Together, these four algorithms—BOSS-FCI, GRaSP-FCI, FCIT, and LV-Dumb—form a continuum.
 The first two establish sound baselines by embedding advanced score-based methods within the GFCI framework.
-FCIT introduces a principled path-based refinement that improves efficiency and reliability, while LV-Lite represents a heuristic extreme: formally incomplete but practically effective.
+FCIT introduces a principled path-based refinement that improves efficiency and reliability, while LV-Dumb represents a heuristic extreme: formally incomplete but practically effective.
 Our empirical evaluation, spanning both simulation and real data, demonstrates the complementary strengths of these approaches.
-BOSS-FCI and GRaSP-FCI retain theoretical correctness but are constrained by repeated testing; FCIT balances correctness with improved precision and scalability; and LV-Lite sacrifices correctness guarantees yet achieves strong accuracy with minimal computational cost in simulation.
+BOSS-FCI and GRaSP-FCI retain theoretical correctness but are constrained by repeated testing; FCIT balances correctness with improved precision and scalability; and LV-Dumb sacrifices correctness guarantees yet achieves strong accuracy with minimal computational cost in simulation.
 The remainder of this paper is organized as follows.
@@
 Section 5 introduces optimization details, Section 6 outlines theoretical guarantees, and Section 8 reports empirical results.
-All algorithms are implemented in the open-source Tetrad suite [12], with reproducibility materials available on GitHub.
+All algorithms are implemented in the open-source Tetrad suite [ramsey2018tetrad], with reproducibility materials available on GitHub.
 Contributions.
@@
   • A targeted-testing algorithm, FCIT, that uses recursive path blocking to maintain well-formed PAGs.
-  • A heuristic, LV-Lite, that outputs the PAG of the BOSS DAG, achieving high empirical accuracy at minimal cost.
+  • A heuristic, LV-Dumb, that outputs the PAG of the BOSS DAG, achieving high empirical accuracy at minimal cost.
   • Empirical validation demonstrating accuracy, scalability, and robustness across simulated linear Gaussian datasets and real datasets.
@@
 We focus on models based on causal directed acyclic graphs (DAGs), which consist of directed edges $X\rightarrow Y$.
-Such an edge indicates that variable $X$ causes variable $Y$, meaning that interventions on $X$ counterfactually alter $Y$ [14].
+Such an edge indicates that variable $X$ causes variable $Y$, meaning that interventions on $X$ counterfactually alter $Y$ [spirtes2001causation].
 A graph $G=(V,E)$ consists of a set of variables $V$ and a set of edges $E$, with at most one edge connecting each vertex pair, where each edge is denoted XYX*{-}*Y.
@@
 An inducing path between $X$ and $Y$ is a path where every non-endpoint node is a collider and an ancestor of $X$ or $Y$. If such a path exists between non-adjacent nodes, the graph is not maximal.
-A MAG or PAG is maximal (following Zhang [17]).
+A MAG or PAG is maximal (following Zhang [zhang2008completeness]).
 if, for every non-adjacent pair $X,Y$, there exists a conditioning set S such that $X$ and $Y$ are m-separated given S.
@@
 A model is edge-minimal of a certain type (e.g., DAG, CPDAG, PAG) if no model of that type has fewer edges while remaining Markov to the data.
-Edge minimality in CPDAGs is discussed by Raskutti and Uhler, [13], Lam et al., [9]; here, we extend this to causally insufficient settings.
+Edge minimality in CPDAGs is discussed by raskutti2018learning, lam2022greedy; here, we extend this to causally insufficient settings.
 A path is blocked by a set $\mathbf{Z}$ if:
@@
 Intuitively, $\text{D-SEP}(A,B)$ contains the variables that may need to be conditioned on to separate $A$ and $B$ when latent confounding is present.
-In the oracle setting, if $A$ and $B$ are not adjacent in the true MAG, then there exists a subset of $\text{D-SEP}(A,B)$ that d-separates them [14].
+In the oracle setting, if $A$ and $B$ are not adjacent in the true MAG, then there exists a subset of $\text{D-SEP}(A,B)$ that d-separates them [spirtes2001causation].
 Given a PAG $G$, the Possible-D-SEP set of node $A$ includes all nodes $V\neq A$ such that there exists a path $U$ between $A$ and $V$ where each triple $\langle X,Y,Z\rangle$ in $U$ satisfies:
@@
   • $Y$ is not a definite non-collider and $\langle X,Y,Z\rangle$ forms a triangle in $G$.
-This set may include nonadjacent nodes and is used by FCI and GFCI to search for separating sets in the presence of latent structure [14].
+This set may include nonadjacent nodes and is used by FCI and GFCI to search for separating sets in the presence of latent structure [spirtes2001causation].
 A discriminating path containing nodes in the path $p=\langle X,...,W,V_{xy},Y\rangle$ in a PAG such that:
@@
 Note that violations of Faithfulness have Lebesgue measure zero if conditional independence is algebraic, though almost unfaithful violations can be common.
-3 Simple Variants of GFCI: BOSS-FCI, GRaSP-FCI, and LV-Lite
-The Greedy FCI (GFCI) algorithm [10] is a hybrid causal discovery method combining score-based search with conditional independence (CI) testing.
+3 Simple Variants of GFCI: BOSS-FCI, GRaSP-FCI, and LV-Dumb
+The Greedy FCI (GFCI) algorithm [ogarrio2016hybrid] is a hybrid causal discovery method combining score-based search with conditional independence (CI) testing.
 It first estimates a Completed Partially Directed Acyclic Graph (CPDAG) using FGES, then refines it into a Partial Ancestral Graph (PAG) through FCI-style edge removal and orientation rules.
@@
 BOSS-FCI.
-BOSS-FCI replaces FGES with the Best Order Score Search (BOSS) algorithm [2].
+BOSS-FCI replaces FGES with the Best Order Score Search (BOSS) algorithm [andrews2019learning].
 BOSS performs permutation-based score search, offering strong accuracy and scalability—especially for dense graphs.
@@
 GRaSP-FCI.
-GRaSP-FCI substitutes the GRaSP algorithm [9] for FGES in the same framework.
+GRaSP-FCI substitutes the GRaSP algorithm [lam2022greedy] for FGES in the same framework.
 GRaSP employs a greedy adjacency search with penalty regularization, emphasizing sparsity and offering a complementary trade-off between accuracy and efficiency.
 Once the CPDAG is obtained, the usual GFCI refinement steps are applied.
-LV-Lite.
-LV-Lite [^3] LV-Lite has been referred to as BOSS-POD (BOSS PAG of DAG) elsewhere, but the name “LV-Lite” has stuck, so we defer.
+LV-Dumb.
+LV-Dumb [^3] LV-Dumb has been referred to as BOSS-POD (BOSS PAG of DAG) elsewhere, but the name “LV-Dumb” has stuck, so we defer.
 is a deliberately simple heuristic.
@@
 The method therefore captures latent-aware patterns implied by the BOSS structure but cannot remove edges to expose bidirected relationships.
-Although formally incomplete, LV-Lite is extremely fast and often achieves empirical accuracy comparable to its more principled counterparts.
+Although formally incomplete, LV-Dumb is extremely fast and often achieves empirical accuracy comparable to its more principled counterparts.
 […]
@@
 Appendix A provides pseudocode for the general *-FCI template, where the initial CPDAG procedure $\mathcal{M}$ may be FGES (standard GFCI), BOSS (BOSS-FCI), or GRaSP (GRaSP-FCI).
-LV-Lite can be viewed as a degenerate instance of this template, omitting edge removal and collider orientation and returning the PAG of the BOSS DAG directly.
+LV-Dumb can be viewed as a degenerate instance of this template, omitting edge removal and collider orientation and returning the PAG of the BOSS DAG directly.
 Input: PAG $\mathcal{G}$ ; CI oracle $\mathcal{T}$
@@
 We now instantiate this FCI-Backward template as the FCIT (Fast Causal Inference Targeted-Testing) algorithm, designed to mitigate the repeated-testing problem in FCI-style searches, including other hybrid score-and-test methods.
-Following GFCI [10], FCIT begins with a score-based procedure that outputs a well-formed CPDAG—e.g., FGES, BOSS, GRaSP, or for small models SP.
+Following GFCI [ogarrio2016hybrid], FCIT begins with a score-based procedure that outputs a well-formed CPDAG—e.g., FGES, BOSS, GRaSP, or for small models SP.
 By default, we use BOSS for its scalability and accuracy.
-As noted by Ogarrio et al., [10], score-based CPDAGs can include extra adjacencies or incorrect orientations when latent variables are present, since a CPDAG cannot represent bidirected edges.
+As noted by ogarrio2016hybrid, score-based CPDAGs can include extra adjacencies or incorrect orientations when latent variables are present, since a CPDAG cannot represent bidirected edges.
 For instance, the MAG $X\to Y\leftrightarrow Z\leftarrow W$ may yield $X\to Y\to Z\leftarrow W$ (with an extra $X\to Z$) or $X\to Y\leftarrow Z\leftarrow W$ (with an extra $W\to Y$).
@@
 The initial CPDAG is converted to a PAG by replacing each edge with $\circ\!$-$\circ$ and transferring the unshielded colliders identified in the CPDAG.
-This is justified because the final PAG must be a subgraph of the CPDAG’s adjacencies, and collider orientations are the only ones that carry over reliably [10].
+This is justified because the final PAG must be a subgraph of the CPDAG’s adjacencies, and collider orientations are the only ones that carry over reliably [ogarrio2016hybrid].
 The refinement phase employs the block_paths_recursively method (Algorithm 4, Appendix A).
@@
 5 Optimizing the Final FCI Orientation Rules
-A key feature of FCIT is that the final FCI orientation rules [17] are applied repeatedly throughout the refinement process rather than only once at the end.
+A key feature of FCIT is that the final FCI orientation rules [zhang2008completeness] are applied repeatedly throughout the refinement process rather than only once at the end.
 This ensures maximally informative orientations but introduces performance bottlenecks, as several rules are computationally expensive in their standard form.
@@
 6.1 Correctness of BOSS-FCI and GRaSP-FCI
-BOSS-FCI and GRaSP-FCI inherit the correctness guarantees of the GFCI algorithm [10].
+BOSS-FCI and GRaSP-FCI inherit the correctness guarantees of the GFCI algorithm [ogarrio2016hybrid].
 In its original form, GFCI included an additional Possible-D-SEP search step that ensures oracle correctness.
@@
 With that step included, BOSS-FCI and GRaSP-FCI are correct for the same reason as GFCI, differing only in the score-based method used for the initial CPDAG.
-6.2 Correctness of LV-Lite
-By contrast, LV-Lite lacks formal correctness guarantees.
+6.2 Correctness of LV-Dumb
+By contrast, LV-Dumb lacks formal correctness guarantees.
 It reports only the PAG equivalence class of the DAG returned by BOSS and therefore cannot remove edges to expose latent confounding or orient bidirected edges.
 Nevertheless, it always produces a structurally well-formed PAG and, as shown in Section 8, performs well on adjacency and orientation accuracy measures.
-To provide intuition for why LV-Lite, while incorrect, nevertheless achieves strong performance on PAG accuracy measures, we note the following considerations.
-It is unsurprising that LV-Lite performs well on arrowhead precision for common adjacencies (AHPC), since it should in principle produce a subset of the arrowheads that exist over the adjacencies shared with the true PAG.
-Similarly, adjacency precision (AP) tends to be high: by GFCI theory, LV-Lite should add only a small fraction of “extra” adjacencies, primarily to cover colliders present in the true PAG.
+To provide intuition for why LV-Dumb, while incorrect, nevertheless achieves strong performance on PAG accuracy measures, we note the following considerations.
+It is unsurprising that LV-Dumb performs well on arrowhead precision for common adjacencies (AHPC), since it should in principle produce a subset of the arrowheads that exist over the adjacencies shared with the true PAG.
+Similarly, adjacency precision (AP) tends to be high: by GFCI theory, LV-Dumb should add only a small fraction of “extra” adjacencies, primarily to cover colliders present in the true PAG.
 Arrowhead recall (AHR) is also expected to be good if the hidden arrowheads represent only a small proportion of those in the true PAG.
-Taken together with the high accuracy of BOSS for producing estimates of CPDAGs, these and other such effects explain why LV-Lite generally yields graphs that are very close to the true PAG—especially in large models—despite its lack of formal guarantees.
+Taken together with the high accuracy of BOSS for producing estimates of CPDAGs, these and other such effects explain why LV-Dumb generally yields graphs that are very close to the true PAG—especially in large models—despite its lack of formal guarantees.
 [^6] We did not pursue, but plan to in future work, estimating effects of direct intervention.
@@
 Proof sketch.
-FCIT uses exactly the arrow- and tail-complete rule set R1–R10 of Zhang, [17].
+FCIT uses exactly the arrow- and tail-complete rule set R1–R10 of zhang2008completeness.
 Given correct adjacencies and unshielded colliders, Spirtes and Zhang prove each rule sound; hence all orientations in $G^{\prime}$ are valid.
@@
 7 Well-formedness of PAGs and Markov
-One observes that FCIT and LV-Lite always return graphs that are well-formed, in the sense that they satisfy strictly the definition of a PAG.
-(For a simulation comparison of how well these various algorithms recover PAGs, consider Figure 12 and related figures in the Appendix.) We can determine this independently by checking whether the identified Zhang MAG calculated from estimated graph [17] satisfies the definition of a MAG and whether when converted back to a PAG the orientations of the estimated PAG are recovered precisely.
+One observes that FCIT and LV-Dumb always return graphs that are well-formed, in the sense that they satisfy strictly the definition of a PAG.
+(For a simulation comparison of how well these various algorithms recover PAGs, consider Figure 12 and related figures in the Appendix.) We can determine this independently by checking whether the identified Zhang MAG calculated from estimated graph [zhang2008completeness] satisfies the definition of a MAG and whether when converted back to a PAG the orientations of the estimated PAG are recovered precisely.
 That is, we check the following
@@
 [^7] The Tetrad suite of software contains such a check.
-By construction, FCIT and LV-Lite always return structurally well-formed PAGs, whereas empirically we find that other FCI-style procedures do not always do so.
+By construction, FCIT and LV-Dumb always return structurally well-formed PAGs, whereas empirically we find that other FCI-style procedures do not always do so.
 In particular:
@@
 Consequently, its output is guaranteed to be structural well-formed regardless of unfaithfulness or finite-sample noise.
-  • LV-Lite.
-LV-Lite trivially returns a well-formed PAG by construction, since it outputs the PAG equivalence class of the BOSS output, which is guaranteed to be a DAG (as BOSS is an algorithm that infers a DAG from a particular permutation of the variables).
+  • LV-Dumb.
+LV-Dumb trivially returns a well-formed PAG by construction, since it outputs the PAG equivalence class of the BOSS output, which is guaranteed to be a DAG (as BOSS is an algorithm that infers a DAG from a particular permutation of the variables).
   • BOSS-FCI GRaSP-FCI, and GFCI. Under unfaithfulness, these algorithms can return malformed graphs in practice.
@@
 Note that each of these algorithms in Oracle mode is guaranteed to produce a well-formed PAG.
-We next compare FCIT with FCI, GFCI, BOSS-FCI, GRaSP-FCI, and the heuristic LV-Lite using simulations implemented in Tetrad [12], focusing on accuracy and efficiency in the presence of latent variables.
+We next compare FCIT with FCI, GFCI, BOSS-FCI, GRaSP-FCI, and the heuristic LV-Dumb using simulations implemented in Tetrad [ramsey2018tetrad], focusing on accuracy and efficiency in the presence of latent variables.
 8 Empirical Evaluation
@@
 [^8] As suggested in an earlier footnote, this is a modeling choice in Tetrad.
-Scores and tests are also available for the multinomial case, the mixed multinomial/Gaussian case, and the mixed multinomial/post-nonlinear case (see [11].
-Also, it should be noted that linear Gaussian scores and tests also work well in the linear non-Gaussian regime [2]).
+Scores and tests are also available for the multinomial case, the mixed multinomial/Gaussian case, and the mixed multinomial/post-nonlinear case (see [ramsey2025scalablecausaldiscoveryrecursive].
+Also, it should be noted that linear Gaussian scores and tests also work well in the linear non-Gaussian regime [andrews2019learning]).
 Tests are also available for the general non-parametric case, though not scores currently; these are less scalable.
@@
 All methods achieve very high adjacency precision.
-FCIT and LV-Lite consistently lead most orientation-precision metrics (AP/AHP, AHPC, arrow-path precision) and always return valid PAGs.
+FCIT and LV-Dumb consistently lead most orientation-precision metrics (AP/AHP, AHPC, arrow-path precision) and always return valid PAGs.
 Tail-path precision is highest for BOSS-FCI and GFCI at large $N$, with FCIT close behind.
-FCIT attains the best bidirected latent-path precision at large samples, while LV-Lite—by design—does not orient bidirected edges.
-In runtime, LV-Lite (and typically BOSS-FCI) are fastest; GFCI is slowest; FCIT and GRaSP-FCI fall in between.
-Results for average degrees 2 and 6 (see Appendix) show qualitatively similar patterns, with denser graphs highlighting the tail-path strengths of BOSS-FCI/GFCI and the robustness of FCIT/LV-Lite on PAG validity.
+FCIT attains the best bidirected latent-path precision at large samples, while LV-Dumb—by design—does not orient bidirected edges.
+In runtime, LV-Dumb (and typically BOSS-FCI) are fastest; GFCI is slowest; FCIT and GRaSP-FCI fall in between.
+Results for average degrees 2 and 6 (see Appendix) show qualitatively similar patterns, with denser graphs highlighting the tail-path strengths of BOSS-FCI/GFCI and the robustness of FCIT/LV-Dumb on PAG validity.
 8.1.4 100 nodes.
 […]
-At this larger scale, LV-Lite and FCIT both maintain exceptionally high arrow-path precision ($\approx 0.99$) and consistently produce structurally valid PAGs, though their tail-path precision is somewhat lower ($\approx 0.93$).
+At this larger scale, LV-Dumb and FCIT both maintain exceptionally high arrow-path precision ($\approx 0.99$) and consistently produce structurally valid PAGs, though their tail-path precision is somewhat lower ($\approx 0.93$).
 In contrast, BOSS-FCI and GRaSP-FCI display the opposite trade-off—achieving very high tail-path precision ($\approx 0.97$) but lower arrow-path precision ($\approx 0.89$).
@@
 Bidirected-edge precision remains low for all methods at this scale, suggesting that in large models such edges should be interpreted cautiously or reserved for qualitative guidance.
-Overall, these results mirror the smaller-graph findings: FCIT offers the best balance of accuracy, scalability, and structural validity, while LV-Lite remains the most efficient and structurally stable baseline.
+Overall, these results mirror the smaller-graph findings: FCIT offers the best balance of accuracy, scalability, and structural validity, while LV-Dumb remains the most efficient and structurally stable baseline.
 8.1.5 Takeaway.
-Across sparsity levels, FCIT and LV-Lite remain the most accurate and stable overall.
+Across sparsity levels, FCIT and LV-Dumb remain the most accurate and stable overall.
 Both consistently produce valid PAGs, sustain near-perfect arrow- and tail-path precision, and recover adjacencies reliably even under latent confounding.
-LV-Lite matches FCIT’s orientation accuracy while running substantially faster, making it a strong baseline when latent structure need not be modeled explicitly.
+LV-Dumb matches FCIT’s orientation accuracy while running substantially faster, making it a strong baseline when latent structure need not be modeled explicitly.
 BOSS-FCI and GRaSP-FCI maintain high directional precision but occasionally yield malformed PAGs and incur heavier computational cost at scale.
@@
 All algorithms maintain very high adjacency precision across sample sizes.
-LV-Lite trails slightly.
-LV-Lite is slightly lower than the others but remains acceptable.
+LV-Dumb trails slightly.
+LV-Dumb is slightly lower than the others but remains acceptable.
 Figure 2: Adjacency Recall (AR) for 20-node graphs with average degree 4.
-LV-Lite and FCIT achieve the highest recall across all sample sizes, reflecting their ability to recover true adjacencies even in the presence of latent confounding.
+LV-Dumb and FCIT achieve the highest recall across all sample sizes, reflecting their ability to recover true adjacencies even in the presence of latent confounding.
 BOSS-FCI, GRaSP-FCI, GFCI, and FCI improve steadily with $N$.
@@
 Figure 3: Arrowhead Precision (AHP) for 20-node graphs with average degree 4.
-LV-Lite and FCIT maintain the highest precision, followed by BOSS-FCI and GRaSP-FCI.
+LV-Dumb and FCIT maintain the highest precision, followed by BOSS-FCI and GRaSP-FCI.
 FCI remains lowest throughout, while GFCI improves moderately at large NN.
@@
 Figure 4: Arrowhead Recall (AHR) for 20-node graphs with average degree 4.
-FCIT and LV-Lite show the strongest growth in recall with increasing NN, approaching 0.6 at large NN.
+FCIT and LV-Dumb show the strongest growth in recall with increasing NN, approaching 0.6 at large NN.
 BOSS-FCI, GFCI, and GRaSP-FCI track closely, while GFCI remains lower overall.
@@
 Figure 5: Arrowhead Precision for Common Adjacencies (AHPC) for 20-node graphs with average degree 4.
-LV-Lite and FCIT achieve the highest arrowhead precision, with LV-Lite reaching 0.92 at large NN.
+LV-Dumb and FCIT achieve the highest arrowhead precision, with LV-Dumb reaching 0.92 at large NN.
 BOSS-FCI and GRaSP-FCI show similar trends, while GFCI and especially FCI remain lower.
@@
 All algorithms improve steadily with increasing NN, with FCI, GRaSP-FCI, and BOSS-FCI converging near 0.9.
-FCIT levels off slightly below these, while LV-Lite stabilizes around 0.85.
+FCIT levels off slightly below these, while LV-Dumb stabilizes around 0.85.
 GFCI performs comparably to the best methods at large samples.
@@
 Figure 7: Tail Precision (TP) for 20-node graphs with average degree 4.
-GFCI, LV-Lite, and LV-Lite lead here, with other algorihtms lagging behind.
+GFCI, LV-Dumb, and LV-Dumb lead here, with other algorihtms lagging behind.
 This metric measures tail precision relative to the true PAG; for a comparison to the directed paths in the true DAG, see Tail Path Precision.
 Figure 8: Tail Recall (TR) for 20-node graphs with average degree 4.
-LV-Lite and FCIT achieve the highest recall across all sample sizes, reflecting their ability to recover correct tail orientations even in the presence of latent confounding.
+LV-Dumb and FCIT achieve the highest recall across all sample sizes, reflecting their ability to recover correct tail orientations even in the presence of latent confounding.
 The other algorithms are grouped substantially lower.
@@
 Figure 9: Arrow Path Precision for 20-node graphs with average degree 4.
-LV-Lite and FCIT again lead in directional precision, followed closely by BOSS-FCI, GRaSP-FCI, and GFCI.
+LV-Dumb and FCIT again lead in directional precision, followed closely by BOSS-FCI, GRaSP-FCI, and GFCI.
 FCI trails across sample sizes.
@@
 Figure 12: Proportion of well-formed PAGs for 20-node graphs with average degree 4.
-As theoretically required, LV-Lite and FCIT achieve perfect PAG validity, whereas other algorithms fall off—sometimes dramatically.
+As theoretically required, LV-Dumb and FCIT achieve perfect PAG validity, whereas other algorithms fall off—sometimes dramatically.
 Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
 All algorithms are quite fast in this regime, with the exception of GFCI, which slows at N=5000N{=}5000.
-LV-Lite lags behind LV-Lite.
+LV-Dumb lags behind LV-Dumb.
 Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
 Algorithm *->-Prec –>-Prec <->-Lat-Prec E-CPU PAG
-LV-Lite 0.9939 0.9320 * 21850.5654 1.0000
+LV-Dumb 0.9939 0.9320 * 21850.5654 1.0000
 FCIT 0.9937 0.9319 0.0000 24430.5061 1.0000
@@
 8.2 Real-Data Analysis
-We analyzed the Algerian Forest Fire dataset from the UCI Machine Learning Repository [1].
+We analyzed the Algerian Forest Fire dataset from the UCI Machine Learning Repository [algerian_forest_fires_547].
 Because we have not discussed nonlinearity or determinism, we removed the six deterministic Canadian Forest Fire System indices—BUI (Build-Up Index), DC (Drought Code), DMC (Duff Moisture Code), FFMC (Fine Fuel Moisture Code), and FWI (Fire Weather Index).
@@
 This left the variables Region (binary), Month, Relative Humidity (RH), Rain, Temperature, Wind Speed (Ws), and Fire (binary occurrence of fire).
-Because the dataset mixes discrete and continuous variables, we used the Degenerate Gaussian score and test [2], with penalty discount $1.0$ and $\alpha=0.05$.
+Because the dataset mixes discrete and continuous variables, we used the Degenerate Gaussian score and test [andrews2019learning], with penalty discount $1.0$ and $\alpha=0.05$.
 This method models discrete variables as multinomial and continuous variables as linearly related, enabling hybrid analysis under a unified likelihood framework.
@@
 Month and Region are correctly identified as exogenous, and the remaining edges are sensible.
-For comparison, LV-Lite returns a PAG nearly identical to the BOSS CPDAG, without introducing bidirected edges. BOSS-FCI and GRaSP-FCI produces similar but slightly denser graphs.
-In Ramsey and Andrews, [11], we conducted a complementary analysis using ChatGPT-4 to provide “expert” judgments of edge and orientation correctness for this dataset.
+For comparison, LV-Dumb returns a PAG nearly identical to the BOSS CPDAG, without introducing bidirected edges. BOSS-FCI and GRaSP-FCI produces similar but slightly denser graphs.
+In ramsey2025scalablecausaldiscoveryrecursive, we conducted a complementary analysis using ChatGPT-4 to provide “expert” judgments of edge and orientation correctness for this dataset.
 This included models where the degenerate Gaussian score was applied without the Canadian Fire System variables (i.e., the same configuration analyzed here), in cases where no domain expert was available.
@@
 FCIT preserved the accuracy of the BOSS initialization while refining orientations in a manner consistent with PAG semantics, highlighting its advantage—when paired with the degenerate Gaussian test and score—for real-world data containing mixtures of continuous and discrete variables.
-Beyond our own benchmark datasets, FCIT and the related nonlinear basis-function test/score [11] have recently been applied to astrophysical data at scale.
-[6] applied FCIT to a highly nonlinear, non-Gaussian sample of roughly $5\times 10^{5}$ low-redshift galaxies from the NASA Sloan Atlas.
+Beyond our own benchmark datasets, FCIT and the related nonlinear basis-function test/score [ramsey2025scalablecausaldiscoveryrecursive] have recently been applied to astrophysical data at scale.
+[desmond2025causalstructuregalacticastrophysics] applied FCIT to a highly nonlinear, non-Gaussian sample of roughly $5\times 10^{5}$ low-redshift galaxies from the NASA Sloan Atlas.
 Using the Basis-Function LRT for independence testing and Basis-Function BIC for scoring, they recovered a causal graph revealing hierarchical, mass-driven relations among galaxy properties while distinguishing them from observational selection effects, which in this note was judged to be quite plausible, recovering known relationships among the studied variables.
@@
 BOSS-FCI and GRaSP-FCI extend the GFCI framework by substituting BOSS or GRaSP for FGES in the initial CPDAG search.
-LV-Lite provides a simple yet effective heuristic alternative that returns the PAG equivalence class of the BOSS DAG, scaling extremely well but lacking the ability to remove edges indicative of latent confounding.
+LV-Dumb provides a simple yet effective heuristic alternative that returns the PAG equivalence class of the BOSS DAG, scaling extremely well but lacking the ability to remove edges indicative of latent confounding.
 Finally, FCIT introduces a targeted-testing strategy that addresses a key statistical limitation of FCI-style methods.
@@
 Empirically, FCIT achieves consistently high adjacency and orientation precision, scales effectively to large models, and outperforms existing FCI-style methods in both accuracy and efficiency.
-Although not formally correct, LV-Lite remains a remarkably fast and accurate heuristic for very large problems.
+Although not formally correct, LV-Dumb remains a remarkably fast and accurate heuristic for very large problems.
 These contributions open several directions for future work.
@@
 Finally, while FCIT infers PAGs that reflect latent confounding, it does not explicitly model the latent variables themselves; integrating latent-variable identification within the same framework remains an important challenge.
-Finally, we would be remiss not to note an important opportunity for future software development relevant to all algorithms in this paper that begin with a BOSS initialization—LV-Lite, BOSS-FCI, and FCIT—for the linear Gaussian or linear non-Gaussian cases.
+Finally, we would be remiss not to note an important opportunity for future software development relevant to all algorithms in this paper that begin with a BOSS initialization—LV-Dumb, BOSS-FCI, and FCIT—for the linear Gaussian or linear non-Gaussian cases.
 Our implementations have focused on Java for portability and integration with the Tetrad platform, which already provides strong performance advantages over many alternatives.
@@
 [^9] The speed improvements are largely due to low-level linear algebra optimizations for calculating likelihoods.
-The practical extension to the linear non-Gaussian case is benchmarked in Andrews et al., [3].
+The practical extension to the linear non-Gaussian case is benchmarked in andrews2023fast.
 For problems with 100 nodes and average degree 10—the “large” case profiled above—CBOSS completes in under a second with performance identical to our Java-based BOSS and scales efficiently to much larger problems.
@@
 In summary, FCIT provides a practical and statistically efficient alternative to existing FCI-style algorithms, consistently yielding well-formed PAGs with improved accuracy and scalability.
-Together with BOSS-FCI, GRaSP-FCI, and the heuristic LV-Lite, it expands the toolkit for causal discovery under latent variables—bridging the gap between correctness, scalability, and practical usability.
+Together with BOSS-FCI, GRaSP-FCI, and the heuristic LV-Dumb, it expands the toolkit for causal discovery under latent variables—bridging the gap between correctness, scalability, and practical usability.
 Acknowledgments and Disclosure of Funding
@@
 This is used in the algorithm, block_paths_recursively, Algorithm 4, to construct sets B on the fly to block all paths between variables $X$ and $Y$.
-This uses the notion of reachability given in [7].
+This uses the notion of reachability given in [geiger1990d].
 The optimization of the discriminating path rule R4 is given in Algorithm 7.
 Section A.2 gives the procedure we use to list discriminating and pre-discriminating paths.
-As an intuition, the recursive blocking procedure underlying FCIT is closely related to the standard depth-first search algorithm for determining d-separation between variables $X$ and $Y$ given a conditioning set $\mathbf{Z}$ [7].
-Its main innovation is to construct the conditioning set dynamically: as the algorithm explores paths from $X$ to $Y$, it incrementally builds a blocking set $\mathbf{B}$ by adding nodes that close open trails—“in a clever way.” [^10] We borrow this turn of phrase from Zhang, [17].
+As an intuition, the recursive blocking procedure underlying FCIT is closely related to the standard depth-first search algorithm for determining d-separation between variables $X$ and $Y$ given a conditioning set $\mathbf{Z}$ [geiger1990d].
+Its main innovation is to construct the conditioning set dynamically: as the algorithm explores paths from $X$ to $Y$, it incrementally builds a blocking set $\mathbf{B}$ by adding nodes that close open trails—“in a clever way.” [^10] We borrow this turn of phrase from zhang2008completeness.
 When the procedure halts and $X$ and $Y$ are not adjacent, it returns a final set $\mathbf{B_{final}}$ such that $\text{d-sep}(X,Y\mid\mathbf{B}_{final})$, or Null if no such separating set exists.
@@
 To certify $m$-separation, both adjacency and path conditions must be satisfied.
-Zhang’s criterion [17] explicitly requires
+Zhang’s criterion [zhang2008completeness] explicitly requires
     $\neg\operatorname{Adj}(x,y,G)\quad\text{and}\quad\forall p\,\operatorname{Blocked}(p,B),$
@@
     C.1 Zhang’s Final FCI Orientation Rules
-    For completeness, we recap here the orientation rules (R0–R10) that Zhang [17] proves are both sound and complete for FCI in the presence of latent variables and selection bias.
+    For completeness, we recap here the orientation rules (R0–R10) that Zhang [zhang2008completeness] proves are both sound and complete for FCI in the presence of latent variables and selection bias.
 In these rules, $\alpha,\beta,\gamma,\theta$ denote generic vertices, and the symbol $\ast$” is a wildcard that can represent an arrowhead ($\rightarrow$), a tail (-), or a circle ($\circ$).
@@
     where each internal collider along the path is a parent of $Y$.
-In the standard approach [14], applying R4 often requires exploring many potential conditioning sets, which is time-consuming and can lead to inaccuracies in dense graphs.
+In the standard approach [spirtes2001causation], applying R4 often requires exploring many potential conditioning sets, which is time-consuming and can lead to inaccuracies in dense graphs.
     To mitigate this, we leverage the recursive blocking procedure (Section 4) to focus only on a small, targeted collection of conditioning sets. Specifically:
@@
 Precision remains uniformly high across all algorithms (>0.93>0.93).
-BOSS-FCI GRaSP-FCI, and GFCI achieve near-perfect accuracy ($\approx 1.0$) throughout, with LV-Lite lagging behimd a bit.
+BOSS-FCI GRaSP-FCI, and GFCI achieve near-perfect accuracy ($\approx 1.0$) throughout, with LV-Dumb lagging behimd a bit.
 FCI is also slightly behind but still quite high.
@@
     Figure 21: Arrowhead Precision (AHP) for 20-node graphs with average degree 2.
-LV-Lite lags at small samples, and FCI is behind overall; the other algorithms do quite well.
+LV-Dumb lags at small samples, and FCI is behind overall; the other algorithms do quite well.
 FCI remains slightly lower but converges steadily.
@@
 Recall improves monotonically with NN for all algorithms, exceeding 0.7 at the largest sample size.
-FCI attains the highest overall recall, followed closely by FCIT, while LV-Lite and the GFCI-family algorithms show similar convergence.
+FCI attains the highest overall recall, followed closely by FCIT, while LV-Dumb and the GFCI-family algorithms show similar convergence.
 Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
     Figure 25: Tail Precision (TP) for 20-node graphs with average degree 2.
-Most algorithms are in a top group, with LV-Lite and FCI lagging behind.
+Most algorithms are in a top group, with LV-Dumb and FCI lagging behind.
 FCI trails slightly at larger sample sizes.
@@
 Precision varies strongly with NN; for large samples.
-LV-Lite is systematically behind.
+LV-Dumb is systematically behind.
 Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
     Figure 30: Proportion of well-formed PAGs for 20-node graphs with average degree 2.
-There is good PAG precision for this regime, with LV-Lite and FCIT returning perfect scores, as predicted.
+There is good PAG precision for this regime, with LV-Dumb and FCIT returning perfect scores, as predicted.
 Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
     Figure 33: Adjacency Recall (AR) for 20-node graphs with average degree 6.
-LV-Lite and FCIT show the strongest gains with sample size, reaching $\approx 0.7$ at large NN.
+LV-Dumb and FCIT show the strongest gains with sample size, reaching $\approx 0.7$ at large NN.
 GFCI, BOSS-FCI, and FCI improve more modestly, and GRaSP-FCI remains lowest.
@@
     Figure 34: Arrowhead Precision (AHP) for 20-node graphs with average degree 6.
-LV-Lite attains the highest precision and improves with NN; FCIT is next best.
+LV-Dumb attains the highest precision and improves with NN; FCIT is next best.
 GFCI and BOSS-FCI remain around $0.7$, GRaSP-FCI is lower, and FCI lowest overall.
@@
     Figure 35: Arrowhead Recall (AHR) for 20-node graphs with average degree 6.
-All curves rise with NN; LV-Lite and FCIT are highest at large NN ($\approx 0.6$), while the GFCI-family methods and FCI trail.
+All curves rise with NN; LV-Dumb and FCIT are highest at large NN ($\approx 0.6$), while the GFCI-family methods and FCI trail.
 Results are averaged over graphs with 0, 4, and 8 latent common causes.
     Figure 36: Arrowhead Precision for Common Adjacencies (AHPC) for 20-node graphs with average degree 6.
-LV-Lite maintains the highest precision and improves with NN; FCIT also increases steadily.
+LV-Dumb maintains the highest precision and improves with NN; FCIT also increases steadily.
 BOSS-FCI and GFCI remain around $0.72$, GRaSP-FCI is lower, and FCI lowest overall.
@@
 Recall increases with NN for all methods.
-At large $N$, BOSS-FCI is highest (near $0.93$), followed by GFCI and FCI; FCIT and GRaSP-FCI are close behind, and LV-Lite is slightly lower.
+At large $N$, BOSS-FCI is highest (near $0.93$), followed by GFCI and FCI; FCIT and GRaSP-FCI are close behind, and LV-Dumb is slightly lower.
 Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
     Figure 39: Tail Recall (TR) for 20-node graphs with average degree 6.
-At the all sample sizes LV-Lite and FCIT have the highest scores, while the others falls short.
+At the all sample sizes LV-Dumb and FCIT have the highest scores, while the others falls short.
 Results are averaged over graphs with 0, 4, and 8 latent common causes.
     Figure 40: Arrow Path Precision for 20-node graphs with average degree 6.
-LV-Lite and FCIT achieve the highest precision, increasing to $\approx 0.96$ at large NN.
+LV-Dumb and FCIT achieve the highest precision, increasing to $\approx 0.96$ at large NN.
 BOSS-FCI and GFCI cluster around $0.80$; GRaSP-FCI is lower and FCI lowest overall.
@@
 FCIT attains the highest precision at mid-range, with other algorithms trailing.
-LV-Lite orients no bidirected edges, yielding undefined precision, so it is not plotted.
+LV-Dumb orients no bidirected edges, yielding undefined precision, so it is not plotted.
 Results are averaged over graphs with 0, 4, and 8 latent common causes.
     Figure 43: Proportion of well-formed PAGs for 20-node graphs with average degree 6.
-LV-Lite and FCIT consistently produce valid PAGs (=1.0=1.0), confirming structural correctness.
+LV-Dumb and FCIT consistently produce valid PAGs (=1.0=1.0), confirming structural correctness.
 Other algorithms fall off, sometimes sharply.
@@
 GRaSP-FCI incurs the highest cost, growing steeply at large NN ($\sim\!600$ ms).
-FCIT is moderate, while LV-Lite, LV-Lite, FCI, and GFCI remain consistently faster.
+FCIT is moderate, while LV-Dumb, LV-Dumb, FCI, and GFCI remain consistently faster.
 Results are averaged over graphs with 0, 4, and 8 latent common causes.